### v1

# Supercritical fluids as a distinct state of matter characterized by sub-short-range structural order

Abstract
A supercritical fluid (SCF) – the state of matter at temperatures and pressures above the critical point – exhibits properties intermediate between those of a liquid and a gas. However, whether it constitutes a fundamentally distinct phase or merely a continuous extension of the liquid and gas states remains an open question. Here we show that a SCF is defined by sub-short-range (SSR) structural order in the spatial arrangement of particles, setting it apart from the gas (disordered), liquid (short-range ordered), and solid (long-range ordered) states. The SSR structural order can be characterized by a length scale effectively quantified by the number of observable peaks in the radial distribution function. This length grows from a minimum microscopic value, on the order of the inter-particle distance at the gas–SCF boundary, to a diverging value at the SCF–liquid boundary. Based on the emergence of SSR order, we demonstrate that the transport and dynamical properties of the SCF state, including the diffusion coefficient, shear viscosity, and velocity autocorrelation function, also clearly distinguish it from both the liquid and gas states. Theoretical predictions are validated by molecular dynamics simulations of argon and further supported by existing experimental evidence. Our study confirms and assigns physical significance to the refined phase diagram of matter in the supercritical region, consisting of three distinct states (gas, supercritical fluid, and liquid) separated by two crossover boundaries that follow universal scaling laws.
Introduction. Liquids and gases are arguably among the most common phases of matter Tabor (1991) and have been studied and probed experimentally since centuries. At low enough pressure and temperature, they are separated by a coexistence line defined by a thermodynamic first order phase transition Domb (2000). Nevertheless, the coexistence line has finite extension and ends into the so called critical point (e.g., $T_{c}\approx 647$ K and $P_{c}\approx 220$ bar for water), above which lies a large part of the phase diagram – the supercritical fluid (SCF) state of matter Kiran et al. (2012); Proctor (2020) – which was first discovered by Baron de la Tour in 1822 de La Tour (1822). In this SCF state, liquid and gas phases can be connected without encountering any thermodynamic singularity, which raises the long-standing question of what really separates a liquid from a gas Brazhkin and Trachenko (2012), if anything.
A commonly considered scenario is that the SCF state is formed by liquid-like and gas-like states, which are separated in the phase diagram by a single separation line (e.g., Widom line, Frenkel line and Widom-Fisher line, etc.) Xu et al. (2005); Jones and Walker (1956); Luo et al. (2014); Banuti et al. (2017); Corradini et al. (2015); Gallo et al. (2014); Fisher and Widom (1969); Vega et al. (1995); Evans et al. (1993); Brazhkin et al. (2012a, 2013); Bolmatov et al. (2013, 2014); Fomin et al. (2018); Nishikawa and Tanaka (1995); Matsugami et al. (2014); Ploetz and Smith (2019); Woodcock (2013, 2016); Brazhkin et al. (2012b); Huang et al. (2023); Ouyang et al. (2024). An alternative scenario is that liquid and solid phases are separated by a crossover region, bounded by two boundary lines Banuti (2015); Maxim et al. (2021); Wang et al. (2021); Li and Jin (2024), making the SCF in this region a distinguishable state from a liquid or a gas. Recent studies have established the thermodynamic origin of these two supercritical boundary lines ($L^{\pm}$ lines)  Li and Jin (2024) and their universal scalings in liquid-gas phase transitions Li and Jin (2024), Ising models Li and Jin (2024), quantum phase transitions Wang et al. (2024); Lv et al. (2024), charged black holes Wang et al. (2025) and liquid-liquid phase transitions Wang et al. (2025).
Figure 1: Phase diagram and radial distribution functions. (a) Phase diagram of argon. The critical point (CP) is located at $(T_{c},P_{c})\approx(151\,\text{K},48.55\,\text{bar})$ Mccain and Ziegler (1967). The liquid-gas coexistence line (solid black line) is obtained from the NIST database NIS . The gray dotted line is the critical isochore, which represents supercritical continuation of the coexistence line. The two black dashed lines are $L^{\pm}$ lines, which are computed using the NIST data and the thermodynamic criterion proposed in Li and Jin (2024), and represent the boundaries between liquid, gas and SCF phases. (b-d) Typical reduced RDF, $\tilde{h}(r)=r[g(r)-1]$, of gas, SCF and liquid states measured in simulations of argon (gas: $T=195$ K, $P=20~\text{bar}$, SCF: $T=195$ K, $90~\text{bar}$, liquid: $T=100$ K, $P=40~\text{bar}$).
If the SCF is an independent and separate state of matter as in the second scenario, complementary to the common liquid, gas and solid states, naturally one would ask what defines this state. Here we address this question focusing on structure, i.e., the spatial arrangement of particles. Our understanding is summarized in Fig. 1 by the phase diagram, and the behavior of a key structural indicator, the radial distribution function (RDF) $g(r)$ that describes the probability of finding a particle at a given distance $r$ from a reference particle. According to textbook knowledge, the structures of gases, liquids and solids (crystals) are characterized by disorder, short-range and long-range order, respectively. These structural features are captured by the behavior of $g(r)$ beyond the first peak, corresponding respectively to, exponential decay, exponentially damped oscillations, and undamped oscillations.
Here we propose that the structure of SCF is characterized by sub-short-range (SSR) order. The SSR order is reflected in $g(r)$ through damped oscillations that are truncated, i.e., the oscillatory behavior of $g(r)$ extends only up to a finite SSR length scale $\xi$. The length scale $\xi$ is ill-defined in the liquid phase, diverges at the liquid-SCF boundary ($L^{+}$ line), and reduces to the scale of nearest neighbor distance at the gas-SCF boundary ($L^{-}$ line).
This viewpoint on structure and, as we will see, also the behavior of transport and dynamical properties, imply that the SCF can be interpreted as an independent state of matter in a loose sense. This definition is not rigorous because there are no bona fide phase transitions between SCF and gases/liquids and all lines should be rather regarded as smooth crossovers. Nevertheless, these three phases can be still separated by two supercritical boundary lines:
(i) Liquid-SCF boundary (LSB). In our framework, the Fisher-Widom line Fisher and Widom (1969) ($\xi\to\infty$) corresponds to the liquid-SCF boundary. In liquids, the oscillations in $g(r)$, $\sim\cos(qr)$, although damped, remain up to $r\to\infty$. The LSB has been alternatively estimated based on dynamic criteria (Frenkel line Brazhkin et al. (2012a)), such as the disappearance of the short-time vibrational motion typical of liquids. The Fisher-Widom line and the Frenkel line are also close to the upper $L^{+}$ line defined from thermodynamic and scaling considerations Li and Jin (2024). In general, the liquid-SCF boundary corresponds to the boundary beyond which the state loses liquid characteristics, although its precise location depends on the detailed criteria. The small quantitative inconsistency is possibly inevitable (also for the gas-SCF boundary), due to the absence of thermodynamic or dynamic singularities. Nevertheless, these lines are close to each other, and can be unambiguously distinguished from the gas-SCF boundary (as we will show in this study). Since the liquid–SCF transition has been extensively studied in the literature, including from the perspective of structural criteria such as the Fisher-Widom line Fisher and Widom (1969); De Carvalho et al. (1994); Vega et al. (1995); Dijkstra and Evans (2000a); Evans et al. (1993); Tarazona et al. (2003); Stopper et al. (2019), we do not focus on it here. Instead, our study concentrates on the gas-SCF boundary, which remains largely unexplored.
(ii) Gas-SCF boundary. This boundary corresponds to the locus where the oscillatory behavior in $g(r)$, aside from the first peak which persists except in the ideal gas limit, disappears. In practice, it can be located by the disappearance of the second peak in $g(r)$, or equivalently the first minimum, as shown by the transition from Fig.1(b) to Fig.1(c), using molecular dynamics (MD) simulation data of argon (Ar). Below, we reveal that this structural change in $g(r)$ is reflected in transport properties, specifically, the viscosity $\eta$ and the self-diffusion coefficient $D$, as well as in the dynamical behavior, such as the velocity autocorrelation function (VACF), $Z(t)$. These observations can also be derived from a liquid-state theory based on the memory-function approach and are supported by simulation results. The numerically estimated gas–SCF boundary based on these criteria aligns closely with the lower $L^{-}$ line independently identified in Ref. Li and Jin (2024) using a thermodynamic criterion, suggesting a clear physical significance for the latter.
Phase diagram and supercritical boundaries. We consider Ar as a prototypical simple system and perform extensive MD simulations (see Methods). In Fig. 1a, we present the phase diagram of Ar. The critical point is located at $(T_{c},P_{c})\approx(151\,\text{K},48.55\,\text{bar})$ Mccain and Ziegler (1967). The liquid-gas coexistence line (the black line in Fig. 1a) continues into the supercritical regime (the gray dotted line), which typically shows a high symmetry revealed by maxima of density fluctuations under constant temperature conditions Xu et al. (2005); Nishikawa and Tanaka (1995); Ploetz and Smith (2019). In this study, we regard all supercritical lines defined in this spirit as the continuation of the coexistence line, including the critical isochore, Widom line Xu et al. (2005), Nishikawa line Nishikawa and Tanaka (1995), and symmetry line Ploetz and Smith (2019). These lines locate slightly differently in the phase diagram (see Li and Jin (2024)), but all around the middle of the SCF phase. Importantly, in our framework, and different from previous interpretation, those are not considered as the “boundary line” between liquid-like and gas-like states.
As shown in Fig. 1a, we propose that the gas, SCF and liquid phases are separated by two boundary lines, (black dashed lines). (i) The liquid-SCF boundary corresponds to the Fisher-Widom line Vega et al. (1995) defined based on the behavior of $g(r)$ (a structural viewpoint), or the Frenkel line defined based on dynamical viewpoints. The location of the Frenkel line depends on the specific criterion employed: the Frenkel-1 line is defined by the equality $\tau_{0}=\tau$ between the vibration time $\tau_{0}$ and the liquid relaxation time $\tau$ Brazhkin et al. (2012a), the Frenkel-2 line corresponds to the disappearance of oscillations and minima of the velocity autocorrelation function Brazhkin et al. (2013), and the Frenkel-3 line is derived based on isochoric heat capacity $C_{V}$ (for monatomic systems such as argon, the criterion is $C_{V}=2k_{\rm B}$Brazhkin et al. (2012a). (ii) The gas-SCF boundary, which is the focus of the present study, has never been individually investigated in previous works. (iii) Several existing considerations have proposed two boundaries, including the $L^{\pm}$ lines Li and Jin (2024), percolation lines Woodcock (2013, 2016), and the two boundaries of the pseudo-boiling transitional region Banuti (2015); Wang et al. (2021). Different criteria give nearby, but quantitatively different boundary lines (see Li and Jin (2024)). Again, the reason is that these lines do not correspond to any thermodynamic or dynamic singularities, which means that they cannot be uniquely defined in a mathematically rigorous way even in the thermodynamic limit. An exception is the $L^{\pm}$ lines in quantum phase transitions, where the thermodynamic supercritical crossovers coincide with dynamic singularities under quantum quenches Wang et al. (2024).
Among the above proposals, only the $L^{\pm}$ lines display universal scalings in various phase transitions Li and Jin (2024); Wang et al. (2024); Lv et al. (2024); Wang et al. (2025): the scalings of the external field and order parameter along $L^{\pm}$ lines satisfy $\delta P^{\pm}\sim(T-T_{\rm c})^{\beta+\gamma}$ and $\delta\rho^{\pm}\sim(T-T_{\rm c})^{\beta}$, where $\beta$ and $\gamma$ are standard critical exponents. Below we will use $L^{\pm}$ lines as a reference of the two boundaries (see Methods).
For the detailed analyses below, we select two isothermal lines corresponding to $T=195$ K (results in the main text) and $T=220$ K (see Supplementary Material) respectively and consider a large range of pressures that scan from largely above the liquid-SCF boundary to well below the gas-SCF boundary. This will allow us to study in detail the structural, transport and dynamical properties across the boundaries, revealing their fundamental origin and nature. In particular, in the next section we will show that the emergence of SSR order in $g(r)$ coincides with the $L^{-}$ line.
Theory. From the Ornstein–Zernike equation Hansen and McDonald (2013), the asymptotic behavior of the reduced total correlation function $\tilde{h}(r)=r[g(r)-1]$ can be expressed in Fourier representation as $\tilde{h}(r)=\sum_{n}A_{n}e^{ik_{n}r}$. The poles $k_{n}$ satisfy the condition $1-\rho\,\hat{c}(k)=0$, where $\rho$ is the number density and $\hat{c}(k)$ is the Fourier transform of the pair direct correlation function $c(r)$. The coefficient $2\pi A_{n}$ represents the residue associated with the $n$-th pole.
In general, an infinite number of poles $k_{n}$ are expected. We emphasize that at least one of these two leading poles must be purely imaginary de Carvalho et al. (1994), while the others consist of complex-conjugate pairs with identical imaginary parts and real parts of opposite sign. However, several studies (e.g., Dijkstra and Evans (2000b)) have shown that retaining only the two leading poles, i.e., those with the smallest imaginary parts, provides an accurate description of $\tilde{h}(r)$ up to interatomic distances of approximately $2\sigma_{0}$, corresponding to second-nearest neighbors ($\sigma_{0}$ is defined by the location of the minimum in the inter-particle potential $V(r)$). Remarkably, this approximation also captures the correct position of the first peak in $\tilde{h}(r)$. This success was attributed to the fact that the next leading poles have a considerably larger imaginary part in a variety of systems de Carvalho et al. (1994); Leote de Carvalho et al. (1999); de Carvalho and Evans (1994); Carvalho and Evans (1997).
Supported by these observations, we then write $\tilde{h}(r)$ as the sum of two dominating terms, $\tilde{h}(r)=A_{0}e^{-r/r_{0}}+A_{1}e^{-r/r_{1}}\cos[q(r-\sigma_{0})]=A_{0}e^{-r/r_{0}}\{1+ae^{-r/\xi}\cos[q(r-\sigma_{0})]\}$, where $a=A_{1}/A_{0}$ and $1/r_{1}=1/r_{0}+1/\xi$ Dijkstra and Evans (2000a). As a direct numerical investigation will confirm, the wave vector $q\approx 2\pi/\sigma_{0}$ is nearly a constant in the interested regime. The first exponential term $e^{-r/r_{0}}$ and the second oscillating term $e^{-r/r_{1}}\cos[q(r-\sigma_{0})]$ correspond respectively to the leading pure imaginary and complex poles. In this language, the Fisher-Widom line can be elegantly defined by the condition $r_{1}=r_{0}$, or equivalently $\xi=\infty$  Fisher and Widom (1969). More precisely, in the liquid phase, $r_{1}>r_{0}$ and the asymptotic decay of $\tilde{h}(r)$ at the limit $r\to\infty$ is dominated by the oscillating term; in the gas phase, $r_{1}<r_{0}$ and the same asymptotic decay is dominated by the exponential term.
Using the two-pole decomposition described above and the newly introduced SSR length scale $\xi$, we propose structural criteria for both the liquid–SCF and gas–SCF boundaries (see Fig. 1(a)). These boundaries were previously identified through thermodynamic analysis Li and Jin (2024), but their precise physical interpretation has remained unclear. Within our proposal, the liquid-SCF boundary is determined by $\xi\to\infty$, which is identical to Fisher-Widom’s condition $r_{1}=r_{0}$ Fisher and Widom (1969); the gas-SCF boundary corresponds to $\xi=c_{0}\sigma_{0}$ where $c_{0}$ is an order one dimensionless constant. From a physical perspective, $\xi$ characterizes the degree of structural correlations in the system. In particular, $\xi=\infty$ corresponds to a strongly correlated many-body system, i.e., a liquid. In contrast, $\xi\sim\mathcal{O}(\sigma_{0})$ indicates a correlation length on the order of the nearest-neighbor distance, as expected in a gas described by a Lennard-Jones potential (Fig. 1(b)), where short-range order is totally lost.
A simple, though approximate, interpretation of $\xi$ is that $n=\xi/\sigma_{0}$ gives the number of peaks observable in $g(r)$. Given the simple expression $1+ae^{-r/\xi}\cos[q(r-\sigma_{0})]$ appearing in $\tilde{h}$, the $n$-th peak at $r=n\sigma_{0}$ is negligible when the second term $ae^{-r/\xi}\cos[q(r-\sigma_{0})]$ is much smaller than one, or $e^{-n\sigma_{0}/\xi}\ll 1$. In this sense, the gas-SCF boundaries correspond to $g(r)$ displaying only a single peak, followed by an exponential decay, as shown in Fig. 1(b). This argument gives a way to approximately estimate $\xi(P)$: as $P$ increases from the dilute limit, when the $n$-the peak appears at $P=P_{n}$, the length scale is evaluated as $\xi(P=P_{n})=n\sigma_{0}$ at $P_{n}$. This idea will guide our data analysis below.
The general picture we expect is the following. (i) In the gas state, the oscillation in $\tilde{h}(r)$ damps rapidly and disappears after the first peak $r\approx\sigma_{0}$ ($\xi/\sigma_{0}\approx n_{0}$; see Fig. 1b). This means that the spatial correlations between particles are negligible beyond a distance comparable to the nearest neighbor distance. The large distance decay of these correlations is determined by the length scale $r_{0}$. (ii) In the SCF state, the SSR structural order gradually grows from the gas-SCF to liquid-SCF boundaries, reflected by the increase of the number of peaks in $\tilde{h}(r)$ from $\xi/\sigma_{0}\approx n_{0}$ to $\xi/\sigma_{0}\to\infty$. In the SCF state, the oscillations are damped and survive up to a finite distance $\xi$ (see Fig. 1c). This corresponds to the gradual formation of higher-order coordination shells around a central tagged particle in the SCF state. (iii) In the liquid state, the oscillations are still damped but extend to an infinitely large distance (see Fig. 1d), as originally advocated by Fisher and Widom Fisher and Widom (1969). The asymptotic decay of these oscillations in the liquid states is characterized by the length scale $r_{1}$.
Although the decomposition in terms of the two length scales $r_{0}$ and $r_{1}$ is mathematically more straightforward, we find that the SSR order length scale, defined as $\xi\equiv r_{0}r_{1}/(r_{0}-r_{1})$, offers a more physically transparent interpretation within the SCF phase. Importantly, all of our conclusions can be equivalently expressed in terms of $r_{0}$ and $r_{1}$, which remain well-defined across all three phases, up to the emergence of long-range order in ideal crystalline states, where they diverge. However, our analysis suggests that reliably extracting $r_{0}$ and $r_{1}$ by directly fitting $g(r)$ requires data of very high precision, making this procedure quite challenging in practice.
Figure 2: SSR structural order and characteristic lengths. (a) Radial distribution function $g(r)$ at different pressures, zoomed in around the second peak, for T=195KT=195\ $\text${K}. (inset) Height of the second peak $h_{2}$ as a function of $P$; below $P_{2}\approx 55$ bar (red circle), the second peak can not be identified anymore. (b) The SSR length $\xi$ and asymptotic decay length $r^{*}$ as functions of $P$ along the isothermal line T=195KT=195\ $\text${K}.
Figure 3: Transport coefficients. (a) Reduced viscosity $\eta^{*}=\eta\left(\sigma^{2}/\sqrt{\epsilon m}\right)$, where $\sigma$ and $\epsilon$ are length and energy parameters in the Lennard-Jones potential (see Methods), and $m$ is the atomic mass. The errors bars represent the standard error of the mean. (b) Reduced diffusion coefficient $D^{*}=D\sqrt{\left(\frac{m}{\epsilon\sigma^{2}}\right)}$. In the kinetic theory expression, $D_{0}=\frac{3}{8\rho\sigma^{2}}\left(\frac{k_{\rm B}T}{\pi m}\right)^{1/2}$, we have used both the ideal gas form $\rho_{\rm ideal}(P,T)=k_{\rm B}P/T$ (ideal kinetic theory), and the exact $\rho(P,T)$ from simulation data (kinetic theory).
Radial distribution functions & physical length scales. We measure $g(r)$ using MD simulations at $T=195$ K, in the pressure range from 5 bar to 900 bar (see Fig. S1a). To provide a characterization of the SSR, we proceed with the simple method introduced above. Let us first focus on the second peak in $g(r)$ (see Fig. 2a). Based on the data of $g(r)$, a separation pressure $P_{2}\approx 55$ bar is determined: when $P>P_{2}$, the height of the second peak $h_{2}(P)$ increases with $P$; when $P<P_{2}$, the second peak cannot be identified within the given accuracy of the numerical data. The position of the second peak, $r_{2}\sim 1.95\sigma_{0}$, provides an estimate of the SSR length, i.e., $\xi(P=P_{2})=r_{2}$. Generalizing this analysis to the third and fourth peaks in $g(r)$, we obtain $\xi(P_{3})=r_{3}$ and $\xi(P_{4})=r_{4}$, where $P_{n}$ is the pressure below which the $n$-th peak in $g(r)$ cannot be identified, and $r_{n}$ is the position of the $n$-th peak (see Fig. S2). Unfortunately, the limitation of the current numerical precision does not allow us to perform such an analysis when $n>4$.
The data of $\xi(P)$, plotted in Fig. 2b, give a direct, though approximate, estimate of the SSR length scale in SCF. Below $P_{2}$, the $g(r)$ has only one peak, $g(r)\approx\exp[-V(r)/k_{\rm B}T]$, as given by the low-density theory Chandler (1987). In this case, $g(r)$ decays beyond the first peak without any oscillations – this is the definition of the gas state. On the other hand, $P_{2}$ should correspond to the gas-SCF boundary at $T=195$ K. Interestingly, $(P_{2}\approx 55\,\rm{bar},T=195\,\rm{K})$ coincides with the point $(P^{-}\approx 60\,\rm{bar},T=195\,\rm{K})$ on the $L^{-}$ line estimated independently in  Li and Jin (2024) using thermodynamics, confirming $L^{-}$ as the gas-SCF boundary. When $P>P_{2}$, a clear growth of the SSR length $\xi(P)$ is observed. If the numerical precision of $g(r)$ is sufficiently high, we expect that $\xi(P)$ would diverge around the liquid-SCF boundary. In sharp contrast, the length scale $r^{*}\equiv\rm{max}(r_{0},r_{1})$, obtained by exponential fitting of the large-$r$ $\tilde{h}(r)$ data (Fig. S3), is nearly a constant in the interested pressure range (Fig. 2b). The results show that $\xi$, which can be unambiguously distinguished from $r^{*}$, captures the onset of SSR in the SCF.
In the Supplementary Material, we verify the universality of these findings by considering a second isothermal cut crossing the $L^{\pm}$ lines (see Fig. S4).
Transport across three states of matter. Next, we present clear signatures of the $L^{\pm}$ crossover lines in the transport coefficients, demonstrating how the emergence of SSR structural order influences their behavior. In the dilute limit, transport processes are well described by the well-known kinetic theory Chapman and Cowling (1990); Hansen and McDonald (2013). However, beyond the gas-SCF boundary, the SSR order provides non-negligible corrections to the transport coefficients - in particular, to the viscosity $\eta$, where the excess part $\eta_{\rm ex}$ that depends on the interaction potential and configurational structure becomes dominant compared to the kinetic part $\eta_{0}$ (see below).
We focus on the diffusion coefficient $D$ and viscosity $\eta$. In MD simulations, $D$ and $\eta$ are computed respectively using the VACF and the stress autocorrelation function, based on the Green-Kubo formula (see Methods for details). For comparison, we also collect the viscosity data from the National Institute of Standards and Technology (NIST) database NIS , which are consistent with our simulation data within numerical accuracy.
To analyze the simulation results and NIST data (see Fig. 3), we first examine the kinetic theory expressions valid in the dilute limit, $D_{0}=\frac{3}{8\rho\sigma_{0}^{2}}\left(\frac{k_{\rm B}T}{\pi m}\right)^{1/2}$ and $\eta_{0}=\frac{5}{16\sigma_{0}^{2}}\left(\frac{mk_{\rm B}T}{\pi}\right)^{1/2}$ Chapman and Cowling (1990); Hansen and McDonald (2013). As shown in Fig. 3a, kinetic theory describes well the viscosity data below the $L^{-}$ line, where $\eta$ is a constant independent of pressure, as expected for gases. However, kinetic theory significantly fails above that line, as $\eta$ exhibits a rapid increase with pressure. The location where this happens is in very good agreement with the thermodynamic estimate of the $L^{-}$ crossover line Li and Jin (2024).
To understand the origin of this deviation, one should note the approximations employed within such a formalism. Kinetic theory, as indicated by its name, only considers the kinetic effects, ignoring any explicit contributions from the interparticle potential. The theory also neglects the memory effects in particle motions by assuming that successive collisions are completely uncorrelated. Mathematically, this means that the memory function of the stress autocorrelation function $G(t)$ is given by $M_{G}(t)\sim\delta(t)$, where $\delta(t)$ is the delta function (see Supplementary Material). As we explicitly demonstrate, both approximations fail above the $L^{-}$ line. Moreover, it is important to emphasize that the microscopic mechanisms underlying fluid viscosity differ fundamentally between gases and liquids. In gases, viscosity arises from particle collisions, whereas in liquids it is governed by potential energy barriers, processes that kinetic theory does not account for. This supports our proposal that the $L^{-}$ line marks the crossover where potential effects become significant and the dynamics are no longer fully described by pair collisions, as in the gaseous regime.
In general, the viscosity contains two terms, $\eta=\eta_{0}+\eta_{\rm ex}$, where $\eta_{0}$ is the kinetic contribution predicted by the kinetic theory, and $\eta_{\rm ex}$ comes purely from the interaction potential. Using this separation, the gas-SCF crossover line can be formally identified as the location at which $\eta_{\rm ex}\approx 0$. Lucks and Moser developed a formalism (hereafter referred to as LM theory) to compute $\eta_{\rm ex}$ based on a memory-function approach Lucas and Moser (1979a, b). In order to achieve a more quantitative description, we take the version of the LM theory that approximates the excess memory function by a simple exponential form $M_{G}(t)=ae^{-bt}$, with $a$ and $b$ explicitly related to $g(r)$ and $V(r)$ (see Supplementary Material). As $P$ increases above $L^{-}$, $\eta_{\rm ex}$ quickly exceeds $\eta_{0}$. Using $g(r)$ from simulations, the theoretical values of $\eta$ computed using the LM theory are very close to simulation and NIST data (see Fig. 3a). Surprisingly, LM theory works reasonably well for both $\eta$ and $D$ even above the $L^{+}$ crossover lines. It remains unclear whether this is merely a fortuitous coincidence. Finally, based on the data of $\eta$, we also define an effective length unit $\sigma_{\rm eff}$ which shows a qualitatively similar behavior of $\sigma_{0}/\sigma_{\rm eff}$ compared to $\xi/\sigma_{0}$ (see Fig. S5). This observation deserves further investigation.
Among transport properties, the diffusion coefficient $D$ presents a more complex case since it is influenced by the structural order in a more indirect way. Because $D$ is related to the integral of the VACF, it does not have an excess term, in addition to the kinetic expression, $D_{0}=\frac{3}{8\rho\sigma_{0}^{2}}\left(\frac{k_{\rm B}T}{\pi m}\right)^{1/2}$. However, because this expression contains the equation of state (EOS) $\rho(P,T)$, which is connected to $g(r)$ and $V(r)$ through the virial equation, the kinetic expression $D_{0}$ depends on structure as well, despite in a more subtle way.
Indeed, $D_{0}$ computed using the full equation of state agrees well with simulation data even beyond the $L^{-}$ line, extending approximately up to the $L^{+}$ line. In contrast, if the ideal gas equation of state is used, i.e., $\rho_{\rm ideal}(P,T)=P/(k_{\rm B}T)$, the resulting “ideal kinetic theory” begins to fail around the $L^{-}$ line (see Fig. 3b). This further supports our interpretation of $L^{-}$ as the boundary where structural correlations beyond nearest neighbors become significant and the system departs from gas-like behavior. Moreover, this analysis highlights that the gas–SCF boundary can be identified through transport coefficients, whose behavior is influenced by the emergence of SSR structural order in the SCF phase.
As already emphasized, above the liquid-SCF boundary (Frenkel-line), transport properties are dominated by a different microscopic mechanism. Meta-stable states in the potential landscape are formed, and the transitions between such states are realized by activation processes. In this regime, the Arrhenius law gives a relation between $D$ and the typical activation energy barrier $\Delta E$, $D\sim e^{-\Delta E/k_{\rm B}T}\sim e^{{\Delta S}/k_{\rm B}}$. The excess entropy $\Delta S$} can be approximately described by the two-particle entropy $S_{2}=-2\pi\rho k_{\rm B}\int_{0}^{\infty}{g(r)\ln[g(r)]-[g(r)-1]}r^{2}dr$ Dzugutov (1996), which can be directly computed from the $g(r)$ data. Leveraging on these arguments, we find that the scaling $D\sim e^{{S_{2}}/k_{\rm B}}$ deviates from the simulation data around $L^{+}$, as expected (see Fig. S6Demmel and Jakse (2025).
Figure 4: Velocity auto-correlation functions. (a) Fitting the tail of the VACF. Assuming a stretched exponential form, $Z(t)\sim e^{-(t/\tau)^{\beta}}$, the exponent $\beta$ is obtained by the slope of the linear fit in the log-log plot of $\ln[Z(t)]$ vs. $t$. (b) $\beta$ as a function of $P$. Error bars represent the standard deviation of the fitted values from ten independent fitting ranges.
Velocity auto-correlation function (VACF). Besides the signatures of the SCF state in structure and transport, the gas-SCF boundary can be also identified by pure dynamic probes, namely the onset of non-trivial memory effects. We focus on the normalized VACF $Z(t)$ (Fig. S1b), which satisfies, $\dot{Z}(t)=-\int_{0}^{t}M_{Z}(t-\tau)Z(\tau)d\tau$, where $M_{z}(t)$ is the memory function. Interestingly, the intermediate-time behavior of the VACF has been used to locate the liquid-SCF boundary (the Frenkel-2 line). In fact, at high pressures, the VACF displays a minimum after a fast decay; this minimum is gradually suppressed by decreasing pressure, which eventually disappears around the crossover pressure defined by the Frenkel-2 line Brazhkin et al. (2013).
Here we propose that the large-time behavior of the VACF, but also the very short one incarnated in the so-called Einstein frequency, contain information that reflects the gas-SCF boundary. It can be shown that (see Supplementary Material), when the memory function is $M_{Z}\sim\delta(t)$ (memory-less) or exponential $M_{Z}(t)\sim e^{-t/\tau_{M}}$, the VACF decays exponentially at large times, $Z(t)\sim e^{-t/\tau}$. However, this simple exponential form is consistent with our simulation data only below the $L^{-}$ line. Above the $L^{-}$ line, we find that at late time (up to the precision of our numerical data) the VACF is well-approximated by a stretched exponential, $Z(t)=A\exp\left[\left(-\frac{t}{\tau}\right)^{\beta}\right]$, with $\beta<1$. In Fig. 4b we show the exponent $\beta$, obtained by fitting the data with the stretched exponential form (Fig. 4a), as a function of pressure. We find that the $L^{-}$ line coincides with the location at which $\beta$ approaches one. The theoretical origin of stretched exponential behavior of $Z(t)$ deserves to be investigated in the future.
In addition, the signature of $L^{\pm}$ lines can also be found in the Einstein frequency $\Omega$ that appears in the very-short-time expansion of $Z(t)=1-\Omega^{2}\frac{t^{2}}{2}+\cdots$ Hansen and McDonald (2013) (see Fig. S7). These results imply that, beyond structural and transport properties, the dynamical behavior of the SCF state also clearly distinguishes it from both the gas and liquid phases, reinforcing its identification as an independent and distinct state of matter.
Figure 5: Experimental diffusion coefficient for supercritical methane at $T=200$ K. Experimental data obtained by quasi-elastic neutron scattering (QENS) Ranieri et al. (2024) and nuclear magnetic resonance (NMR) Oosting and Trappeniers (1971) are shown respectively with open circles and open diamonds. The green line is the theoretical prediction from ideal kinetic theory, with scaling $D^{*}\propto P^{-1}$.
Experimental signatures and proposals. Our findings provide a basis for experimental verification in future studies. Accurate measurements of $g(r)$ in the SCF state are becoming available using modern neutron scattering instruments Dove and Li (2022); Iwase et al. (2010) and the state-of-the-art data analysis tools Ma et al. (2020). This would allow one to directly extract the SSR length $\xi$ from the experimental data using the approach presented here. It is also possible to measure transport coefficients using neutron scattering techniques.
In fact, a recent experimental study has already reported accurate data of $D$ in supercritical methane along the $T=200$ K isotherm Ranieri et al. (2024). In Fig.5, we present the experimental data from Refs. Oosting and Trappeniers (1971); Ranieri et al. (2024), together with the $L^{\pm}$ lines determined based on the NIST EOS of methane. The experimental data are fully consistent with our theoretical expectation. The diffusion constant $D$ follows the ideal kinetic theory behavior $D\sim 1/P$ up to a pressure around the $L^{-}$ line as in Fig. 3b. Beyond the $L^{-}$ line a clear deviation from this scaling is observed. The SCF state between $L^{\pm}$ lines appears to coincide with a “transition regime” between the gaseous low-$P$ and the liquid high-$P$ behavior. In agreement with our simulation results, this observation provides a first experimental evidence of the two $L^{\pm}$ crossover lines and of the existence of the supercritical fluid state, clearly separated from the gas and liquid phases.
Finally, the signatures of the supercritical boundary lines could be experimentally observed in a more phenomenological way. Enhanced density fluctuations have been observed in small-angle neutron scattering experiments around the $L^{+}$ line Li and Jin (2024) in supercritical carbon dioxide Pipich and Schwahn (2018), due to the formation of liquid-like droplets. Analogously, we also expect enhanced density fluctuations around the $L^{-}$ line, due to formation of gas-like bubbles. If one interprets the supercritical transition regime between $L^{\pm}$ lines as the blur-up of the subcritical discontinuous first-order transition, then the rapid change of the EOS in this regime might be related to the phenomenon of supercritical pseudo-boiling Banuti (2015); Wang et al. (2021); He et al. (2023).
Discussion. In this study, we demonstrate that the signatures of the boundary between gas and SCF states can be identified by various physical quantities, including structure descriptors such as the radial distribution function, transport coefficients, and the velocity auto-correlation function. There have been extensive studies on the other boundary, i.e., the liquid-SCF boundary. Our study completes the phase diagram above the critical point clarifying the existence of three independent states of matter characterized by distinct physical properties. (i) The liquid-SCF boundary is the limit beyond which the system does not behave like a typical liquid. Because the liquid state can be characterized by various features, ranging from thermodynamics to dynamics to structures, the deviation from the liquid behavior can be defined according to different criteria Fisher and Widom (1969); Brazhkin et al. (2012a, 2013, a). (ii) The gas-SCF boundary is the limit beyond which the system does not behave like a typical gas. This line is reflected in several physical properties and can be therefore defined using thermodynamics, structural order, transport coefficients and finally dynamical behavior. (iii) In-between is the SCF state that is neither liquid-like nor gas-like but rather a separate state of matter, whose properties cannot be easily rationalized as a simple linear superposition of gas and liquid features.
It remains unclear whether our findings, and in particular the existence of a gas-SCF boundary, extend to systems with purely repulsive interactions, where no thermodynamic separation between liquid and gas phases, and no critical point, exist. So far, in these systems, only a boundary line between gas-like and liquid-like states has been discussed, e.g., Huang et al. (2023); Jiang et al. (2025).
On the other hand, the proposed framework and analysis methods can be generalized to many other systems. In particular, it would be very interesting to study the nature of order in the quantum supercritical state Wang et al. (2024); Lv et al. (2024), in the context of quantum many-body correlations. Moreover, the dynamical signatures of the $L^{-}$ line can be explored by in the supercritical kinetics of black holes Zhao et al. (2025a); Li et al. (2025a, b) and holographic models as well Zhao et al. (2025b). Another future direction is to examine in more detail the analogy between the supercritical fluid state and the QCD phase diagram of nuclear matter. Interestingly, a three-state phase diagram presenting an intermediate partially deconfined phase, and naive similarities with our proposal depicted in Fig. 1(a), has been recently discussed Hanada and Watanabe (2023); Glozman et al. (2022); Fujimoto et al. (2025).
In conclusion, returning to the fundamental question of “what separates a liquid from a gas” Brazhkin and Trachenko (2012), our revised answer points to the existence of a distinct supercritical fluid state, clearly distinguishable from both gas and liquid phases in terms of thermodynamic, structural, transport, and dynamical properties.
Methods
Molecular dynamics simulations
In MD simulations, we employ the Lennard-Jones (LJ) potential for argon in the LAMMPS package Thompson et al. (2022), which benefits from efficient parallelization for large atom counts. The expression for the LJ potential is given as
$V(r)=4\epsilon\left[\left(\frac{\sigma}{r}\right)^{12}-\left(\frac{\sigma}{r}\right)^{6}\right],$ (1)
where $\epsilon/k_{\rm B}=119.5$ K, $\sigma=$3.405\text{\,}\mathrm{\text{\AA }}$\$, and $k_{\rm B}=$1.38\text{\times}{10}^{-23}\text{\,}\mathrm{J}\text{\,}{\mathrm{K}}^{-1}$\$ is the Boltzmann constant. The initial system is a face-centered cubic lattice of 4000 atoms with a lattice constant of $5.26\text{\,}\mathrm{\text{\AA }}$. Thirty-five different pressures were simulated from $5\text{\,}\mathrm{bar}$ to $1300\text{\,}\mathrm{bar}$. For each target pressure, the system is initially relaxed at $195\text{\,}\mathrm{K}$ for $5\text{\,}\mathrm{ns}$ in an isothermal-isobaric (NPT) ensemble, in order to reach thermal equilibrium. Then the simulation is performed in a canonical (NVT) ensemble for $20\text{\,}\mathrm{ns}$ – all measurements are performed in this time period. Temperature and pressure are regulated via the Nosé-Hoover methods Nosé (1984); Hoover (1985), with Newton’s equations of motion solved numerically using a time step of $1\text{\,}\mathrm{fs}$. All samples are simulated with periodic boundary conditions. The pressure unit is bar, length unit Angstroms, time unit fs, and energy unit kcal/mol.
The diffusion coefficient $D$ is calculated using the Einstein relation from the mean square displacement (MSD) of a 15 ns NVT equilibrium trajectory at each pressure. The viscosity $\eta$ is calculated using the Green–Kubo formalism, which relates $\eta$ to the integral of the stress autocorrelation function. Specifically, $\eta$ is given by
$\begin{split}\eta=\frac{1}{Vk_{\rm B}T}\int_{0}^{\infty}\left\langle S^{\alpha\beta}(0)S^{\alpha\beta}(t)\right\rangle dt,\\ S^{\alpha\beta}=-\sum_{i}^{N}\left(mv_{i}^{\alpha}v_{i}^{\beta}+f_{i}^{\alpha}r_{i}^{\beta}\right),\end{split}$ (2)
where $S^{\alpha\beta}(t)$ ($\alpha\neq\beta$) denotes an off-diagonal component of the microscopic stress tensor, expressed as the sum of kinetic ($mv_{i}^{\alpha}v_{i}^{\beta}$) and virial ($f_{i}^{\alpha}r_{i}^{\beta}$) contributions. In our simulations, the stress autocorrelation function is computed from a long-time NVT trajectory of at least 15 ns. The time correlation is averaged over multiple time origins along the trajectory, ensuring the convergence of the Green–Kubo integral within the correlation time. For low-pressure systems, longer trajectories are used to capture the slower decay of stress fluctuations, and the average over multiple independent simulations with different initial conditions are performed to further reduce the statistical noise.
Thermodynamic definition of $L^{\pm}$ lines
To define from thermodynamics the $L^{\pm}$ lines Li and Jin (2024), one first chooses the critical isochore, which is an extension of the coexistence line to the supercritical region, as the reference line. The compressibility $\kappa_{T}\equiv\frac{P_{c}}{\rho^{\rm m}_{c}}\left(\frac{\partial\rho^{\rm m}}{\partial P}\right)_{T}$ is evaluated along each path parallel to the critical isochore, where $P_{c}$ and $\rho^{\rm m}_{c}$ are the critical pressure and critical mass density. Since $\kappa_{T}$ is a function of distance $\delta P(\rho^{\rm m},T)=P(\rho^{\rm m},T)-P(\rho^{\rm m}_{c},T)$ and $T$, one can find a temperature $T_{\rm max}(\delta P)$ that maximizes $\kappa_{T}$ along each path with a fixed $\delta P$. All the $T_{\rm max}(\delta P)$ points under different $\delta P$ together consist of $L^{\pm}$ lines. In order to determine $L^{\pm}$ lines, one needs to know the EOS $P(\rho^{\rm m},T)$. In this study, we take the EOS of argon from the NIST database NIS . The resulting $L^{\pm}$ lines are plotted in Fig. 1a.
Acknowledgements.
SJ and MB acknowledge the support of the Shanghai Municipal Science and Technology Major Project (Grant No.2019SHZDZX01). XL acknowledges the Postdoctoral Fellowship Program of CPSF (Grant Number GZC20252776). XF acknowledges the support of the National Natural Science Foundation of China (No.52201016). MB acknowledges the support of the sponsorship from the Yangyang Development Fund. YJ acknowledges funding from NSFC (Grant 12047503) and Wenzhou Institute (Grant WIUCASICTP2022). SJ and XF thank the Key State Laboratory for Clean and Efficient Turbomachinery Power Equipment for providing HPC resources. The authors acknowledge the use of the High Performance Cluster at Institute of Theoretical Physics, Chinese Academy of Sciences.
Supplementary Material
In the supplementary material we provide extended data, details of the analyses presented in the main text and additional technical information about the theoretical frameworks used.
Figure S1: Extended data. (a) Radial distribution functions. Data for argon are obtained from MD simulations at P=5,15,25,35,45,55,60,75,90,105,120,135,150,200,320,500,700,800barP=5,15,25,35,45,55,60,75,90,105,120,135,150,200,320,500,700,800\ $\text${bar} (from red to blue) and T=195KT=195\ $\text${K}. (b) Velocity auto-correlation functions. Data for argon are obtained from MD simulations at P=5,10,15,20,25,35,45,55,60,75,90,100,110,120,135,150,200,210,320,400,500,600,700,800,900,1000,1100,1200,1300barP=5,10,15,20,25,35,45,55,60,75,90,100,110,120,135,150,200,210,320,400,500,600,700,800,900,1000,1100,1200,1300\ $\text${bar} (from right to left) and T=195KT=195\ $\text${K}.
Figure S2: Estimate of the SSR length $\xi$. Radial distribution functions around the (a) third ($n=3$) and (b) fourth ($n=4$) peaks. (inset) The height of the third ($h_{3}$) and fourth ($h_{4}$) peaks as a function of $P$.
Figure S3: Evaluation of the length scale $r^{*}$. Here $r^{*}$ is obtained by an exponential fit of the large-$r$ data of $\tilde{h}(r)=r[g(r)-1]$. (a) For $P<P^{+}=285$ bar (below the $L^{+}$ line), the fitting is performed directly to $\tilde{h}(r)$. (b) For $P>P^{\rm+}$ (above the $L^{+}$ line), the fitting is performed to the envelope of peaks. Note that when $P<P^{+}$, there are only a few peaks at small $r$; the tail of $\tilde{h}(r)$ does not show peaks.
Figure S4: SSR structural order and characteristic lengths along the isothermal line T=220KT=220\ $\text${K}. Red line and blue line represent the SSR length $\xi$ and asymptotic decay length $r^{*}\equiv\max(r_{0},r_{1})$ as functions of $P$ respectively.
Figure S5: Effective length unit from the viscosity data. We assume that the kinetic expression of viscosity can be generalized to higher pressures, with a variable effective particle diameter $\sigma_{\rm eff}(P)$, $\eta=\frac{5}{16\sigma_{\rm eff}^{2}(P)}\left(\frac{mk_{\rm B}T}{\pi}\right)^{1/2}$, where $\sigma_{\rm eff}$ is obtained by fitting the NIST data of $\eta$. The growth of $\sigma_{0}/\sigma_{\rm eff}$ begins around the $L^{-}$ line.
Figure S6: Reduced diffusion coefficient as a function of the two-particle structural entropy $S_{2}$. The solid line represents an exponential fit of the data above the $L^{+}$ line, $D=1.18e^{S_{2}/k_{\rm B}}$.
Figure S7: Einstein frequency $\Omega$ (Thz) as a function of pressure $P$. The power-law fits (lines) are guides to the eye.
Theories of auto-correlation functions and transport coefficients
Velocity auto-correlation function (VACF) and diffusion coefficient
The (normalized) VACF is defined by,
$Z(t)\equiv\frac{\left\langle{\bf v}(0)\cdot{\bf v}(t)\right\rangle}{\left\langle{\bf v}(0)\cdot{\bf v}(0)\right\rangle},$ (S1)
and the diffusion coefficient $D$ is related to the VACF through the Green-Kubo formula,
$D=\frac{k_{\rm B}T}{m}\int_{0}^{\infty}Z(t)dt,$ (S2)
where $T$ is the temperature, $m$ the mass of particle, $k_{\rm B}$ the Boltzmann constant. In general, the VACF satisfies the following equation,
$\dot{Z}(t)=-\int_{0}^{t}M_{Z}(t-\tau)Z(\tau)d\tau,$ (S3)
where $M_{Z}(t)$ is the memory function describing the memory effects in the motion of particles, and $\dot{Z}$ represents the time derivative of $Z$.
Kinetic theory
In the dilute limit, the dynamics are memory-less,
$M_{Z}(t)=\frac{1}{\tau}\delta(t),$ (S4)
and the VACF decays exponentially,
$Z(t)=e^{-t/\tau},$ (S5)
where $\tau$ is the velocity correlation time. Equation (S2) then gives,
$D=\frac{k_{\rm B}T\tau}{m}.$ (S6)
Elementary kinetic theory Chapman and Cowling (1990) shows that in the dilute limit, $\tau=l/\langle v\rangle$, where $l$ is the mean free path and $\langle v\rangle$ the average velocity. For hard spheres, the kinetic theory gives, $\tau=\frac{3}{8\rho\sigma_{0}^{2}}\left(\frac{\pi k_{\rm B}T}{m}\right)^{-1/2}$, where $\sigma_{0}$ is the diameter of hard spheres, and
$D_{0}=\frac{3}{8\rho\sigma_{0}^{2}}\left(\frac{k_{\rm B}T}{\pi m}\right)^{1/2}.$ (S7)
For the Lennard-Jones potential, $r=\sigma_{0}\approx 1.12\sigma$ is where the potential energy $V(r)$ (see Eq. 1) minimizes.
Lucks-Moser theory
At high densities, the memory-less assumption, Eq. (S4), does not hold anymore. However, in the interested regime, i.e., for the SCF around the gas-SCF boundary, the memory effects are not significant, and the memory function $M_{Z}(t)$ decays much faster than the VACF $Z(t)$. In this case, one can assume a simple form of $M_{Z}(t)$, such as the following exponential form,
$M_{Z}=\Omega^{2}e^{-t/\tau_{M}}.$ (S8)
Here $\Omega$ is the Einstein frequency that appears in the short-time expansion of $Z(t)=1-\Omega^{2}\frac{t^{2}}{2}+\cdots$, and $\tau_{M}$ is the characteristic time of the memory function. Plugging Eq. (S8) into into Eq. (S3), one obtains,
$Z(t)=\frac{1}{\alpha_{+}-\alpha_{-}}\left(\alpha_{+}e^{-\alpha_{-}t}-\alpha_{-}e^{-\alpha_{+}t}\right),$ (S9)
where
$\alpha_{\pm}=\frac{1}{2\tau_{M}}\left[1\mp\left(1-4\Omega^{2}\tau_{M}^{2}\right)^{1/2}\right].$ (S10)
Equation (S2) then gives,
$D=\frac{k_{\rm B}T}{m\Omega^{2}\tau_{M}}.$ (S11)
In order to explicitly compute $D$, one needs to know $\Omega^{2}$ and $\tau_{M}$. Lucks and Moser has related these two quantities to $g(r)$ and the inter-particle potential $V(r)$, through the following formulas Lucas and Moser (1979a):
$\Omega^{2}=\frac{\left\langle\dot{\bf v}(0)\cdot\dot{\bf v}(0)\right\rangle}{\left\langle{\bf v}(0)\cdot{\bf v}(0)\right\rangle},$ (S12)
$\frac{1}{\tau_{M}}=\left[\frac{\left\langle\ddot{\bf v}(0)\cdot\ddot{\bf v}(0)\right\rangle}{\left\langle\dot{\bf v}(0)\cdot\dot{\bf v}(0)\right\rangle}-\frac{\left\langle\dot{\bf v}(0)\cdot\dot{\bf v}(0)\right\rangle}{\left\langle{\bf v}(0)\cdot{\bf v}(0)\right\rangle}\right]^{1/2},$ (S13)
where
$\left\langle{\bf v}(0)\cdot{\bf v}(0)\right\rangle=\frac{3k_{\rm B}T}{m},$ (S14)
$\frac{\left\langle\dot{\bf v}(0)\cdot\dot{\bf v}(0)\right\rangle}{3k_{\rm B}T/m\sigma_{0}}=32\pi\frac{n^{*}}{T^{*}}\int_{0}^{\infty}g_{2}\left(r_{12}^{*}\right)\left[3W_{12}^{\prime}+2W_{12}^{\prime\prime}r_{12}^{*2}\right]r_{12}^{*2}dr_{12}^{*},$ (S15)
[…]
with
$\begin{aligned} W_{ij}=\frac{V(r)}{4\epsilon},\quad W_{ij}^{\prime}=dW_{ij}/dr_{ij}{}^{*2},\quad W_{ij}^{\prime\prime}=d^{2}W_{ij}/\left(dr_{ij}{}^{*2}\right)^{2},\quad n^{*}=n\sigma_{0}^{3},\quad T^{*}=\frac{k_{\rm B}T}{\epsilon}.\end{aligned}$ (S17)
Equations (S11)-(LABEL:eq:ddv2) give a theoretical estimate of the diffusion coefficient, $D_{\rm LM}$. Compared to the kinetic theory $D_{0}$, $D_{\rm LM}$ takes into account the dynamical memory effects (Eq. S8) and the spatial pair correlations $g(r)$.
Unsurprisingly, even the Lucks-Moser theory breaks down if the density is further increased. The VACF Eq. (S9) still has an exponential form at large times, with a correlation time $\tau\sim 1/\alpha_{-}$ (note that $\alpha_{+}<\alpha_{-}$ ). As shown in the main text, this exponential form does not hold anymore at high densities – instead, we need a stretched exponential function to fit the simulation data. However, the exponential approximation does not seem to affect significantly the transport coefficients.
Stress auto-correlation function and Viscosity
The viscosity is connected to the stress auto-correlation function,
$G(t)\equiv\left\langle S^{\alpha\beta}(0)S^{\alpha\beta}(t)\right\rangle,$ (S18)
via the Green-Kubo relation,
$\eta=\frac{1}{Vk_{\rm B}T}\int_{0}^{\infty}G(t)dt.$ (S19)
Here $S^{\alpha\beta}$ represents the $\alpha\beta$ component of the off-diagonal ($\alpha\neq\beta$) stress tensor, defined as,
$S^{\alpha\beta}=-\sum_{i}^{N}\left(mv_{i}^{\alpha}v_{i}^{\beta}+f_{i}^{\alpha}r_{i}^{\beta}\right)=S^{\alpha\beta}_{0}+S_{\rm ex}^{\alpha\beta},$ (S20)
where $f_{i}^{\alpha}$ is the net force of all particles on the $i$-th particle in the $\alpha$-direction, and $r_{i}^{\beta}$ is the coordinate of the $i$-th particle in the $\beta$-direction. For the convenience of discussion below, we separates the kinetic stress $S_{0}^{\alpha\beta}$ that depends on velocities, and the excess stress $S_{\rm ex}^{\alpha\beta}$ that depends on the forces (the potential energy). Because the considered fluid is isotropic, we assume, $S^{xy}=S^{yz}=S^{xz}$.
Kinetic theory
Using Eqs. (S18) and (S20), we can expand $G(t)$ into four terms,
$\begin{split}G(t)&=\left\langle S_{0}^{\alpha\beta}(0)S_{0}^{\alpha\beta}(t)\right\rangle+\left\langle S_{\rm ex}^{\alpha\beta}(0)S_{\rm ex}^{\alpha\beta}(t)\right\rangle+\left\langle S_{0}^{\alpha\beta}(0)S_{\rm ex}^{\alpha\beta}(t)\right\rangle+\left\langle S_{\rm ex}^{\alpha\beta}(0)S_{0}^{\alpha\beta}(t)\right\rangle\\ &\approx\left\langle S_{0}^{\alpha\beta}(0)S_{0}^{\alpha\beta}(t)\right\rangle+\left\langle S_{\rm ex}^{\alpha\beta}(0)S_{\rm ex}^{\alpha\beta}(t)\right\rangle\\ &\equiv G_{0}(t)+G_{\rm ex}(t).\end{split}$ (S21)
The third and fourth terms in the first line are neglected, assuming that the kinetic and potential parts are uncorrelated. In the kinetic theory, the excess part $G_{\rm ex}(t)$ is also ignored. In this case, $\eta$ is determined solely by velocity auto-correlations. Under similar approximations (e.g., the dynamics are memoryless), one can derive Chapman and Cowling (1990),
$\eta_{0}=\frac{5}{16\rho\sigma_{0}^{2}}\left(\frac{mk_{\rm B}T}{\pi}\right)^{1/2},$ (S22)
which has a simple relation with $D_{0}$, $\eta_{0}\sim D_{0}\rho$.
Lucks-Moser theory
Lucks and Moser have also computed the excess viscosity by considering an exponential memory function Lucas and Moser (1979b). For the excess part, the stress auto-correlation function is,
$G_{\rm ex}(t)\equiv\left\langle S_{\rm ex}^{\alpha\beta}(0)S_{\rm ex}^{\alpha\beta}(t)\right\rangle,$ (S23)
which satisfies,
$\dot{G}_{\rm ex}(t)=-\int_{0}^{t}M_{G}(t-\tau)G_{\rm ex}(\tau)d\tau.$ (S24)
Using an exponential ansatz,
$M_{G}(t)=ae^{-bt},$ (S25)
where
$a=\frac{\left\langle\dot{S}^{\alpha\beta}(0)\dot{S}^{\alpha\beta}(0)\right\rangle}{\left\langle{S}^{\alpha\beta}(0){S}^{\alpha\beta}(0)\right\rangle},$ (S26)
$b=\left[\frac{\left\langle\ddot{S}^{\alpha\beta}(0)\ddot{S}^{\alpha\beta}(0)\right\rangle}{\left\langle\dot{S}^{\alpha\beta}(0)\dot{S}^{\alpha\beta}(0)\right\rangle}-\frac{\left\langle\dot{S}^{\alpha\beta}(0)\dot{S}^{\alpha\beta}(0)\right\rangle}{\left\langle{S}^{\alpha\beta}(0){S}^{\alpha\beta}(0)\right\rangle}\right]^{1/2},$ (S27)
the excess viscosity can be evaluated,
$\eta_{\rm ex}=\frac{b}{Vk_{\rm B}Ta},$ (S28)
and the total viscosity is,
$\eta=\eta_{0}+\eta_{\rm ex}.$ (S29)
Note that the $\eta_{\rm ex}$ depends $g(r)$ through the following expressions,
$\frac{\left\langle{S}^{\alpha\beta}(0){S}^{\alpha\beta}(0)\right\rangle}{V(kT)^{2}\left(1/\sigma_{0}^{3}\right)}=n^{*}+\frac{16\pi}{15}\frac{n^{*2}}{T^{*}}\int_{0}^{\infty}g_{2}\left(r_{12}^{*}\right)\left[5W_{12}^{\prime}+2W_{12}^{\prime\prime}r_{12}^{*2}\right]r_{12}^{*4}dr_{12}^{*},$ (S30)
$\begin{aligned} &\frac{\left\langle\dot{S}^{\alpha\beta}(0)\dot{S}^{\alpha\beta}(0)\right\rangle}{V(kT)^{2}\left(\epsilon/m\sigma_{0}^{5}\right)}=\frac{1024\pi}{15}\frac{n^{*2}}{T^{*}}\int_{0}^{\infty}g_{2}\left(r_{12}^{*}\right)\left[10W_{12}^{\prime}+4W_{12}^{\prime}W_{12}^{\prime\prime}r_{12}^{*2}+W_{12}^{\prime\prime}r_{12}^{*4}\right]r_{12}^{*4}dr_{12}^{*},\end{aligned}$ (S31)
$\begin{aligned} \frac{\left\langle\ddot{S}^{\alpha\beta}(0)\ddot{S}^{\alpha\beta}(0)\right\rangle}{V(kT)^{2}\left(\epsilon^{2}/m^{2}\sigma_{0}^{7}\right)}=&\frac{2048\pi}{15}\frac{n^{*2}}{T^{*}}\int_{0}^{\infty}g_{2}\left(r_{12}{}^{*}\right)\left[T^{*}\left(90W_{12}^{\prime}+180W_{12}^{\prime}W_{12}^{\prime\prime}r_{12}{}^{*2}\right.\right.+24W_{12}^{\prime}W^{\prime\prime\prime}{}_{12}r_{12}{}^{*4}\\ &+231W^{\prime\prime}{}_{12}{}^{2}r_{12}{}^{*4}+84W^{\prime\prime}{}_{12}W^{\prime\prime\prime}{}_{12}r_{12}{}^{*6}\left.+12W^{\prime\prime\prime}{}_{12}{}^{2}r_{12}{}^{*8}\right)+4\left(20W_{12}^{\prime}+24W_{12}^{\prime}W^{\prime\prime}{}_{12}r_{12}{}^{*2}\right.\\ &\left.\left.+18W_{12}^{\prime}W^{\prime\prime}{}_{12}{}^{2}r_{12}{}^{*4}+4W^{\prime\prime}{}_{12}{}^{3}r_{12}{}^{*6}\right)r_{12}{}^{*2}\right]r_{12}{}^{*2}dr_{12}{}^{*},\end{aligned}$ (S32)
where
$\begin{aligned} W_{ij}^{\prime\prime\prime}&=d^{3}W_{ij}/\left(dr_{ij}{}^{*2}\right)^{3}.\end{aligned}$ (S33)