### v2

# Physical Signatures of Supercritical Fluid Boundaries

Abstract
In the supercritical fluid (SCF) region, at temperatures and pressures above the critical point, the thermodynamic singularity separating liquids and gases no longer exists. Recent arguments based on thermodynamics and critical scalings have revived the proposal that the SCF constitutes an intermediate state of matter, separated from the liquid and gas by two supercritical boundaries, the $L^{\pm}$ lines. However, until now, the nature of the supercritical state and the physical signatures of these boundaries have remained elusive. Here, we demonstrate that the SCF is characterized by distinct structural, transport, and dynamical behavior. Specifically, the spatial arrangement of particles—captured by the radial distribution function—as well as the diffusion coefficient, shear viscosity, and velocity autocorrelation function in the SCF regime are qualitatively different from those in both the liquid and gas states and exhibit clear physical signatures upon crossing the $L^{\pm}$ lines. Our theoretical predictions are validated by molecular dynamics simulations of argon and are further supported by existing experimental evidence. These results provide a clear physical foundation for a refined phase diagram of matter in the supercritical region, comprising three distinct states—gas, supercritical fluid, and liquid—separated by two crossover boundaries obeying universal scaling laws.
Introduction. Liquids and gases are arguably among the most common phases of matter Tabor (1991) and have been studied and probed experimentally since centuries. At low enough pressure and temperature, they are separated by a coexistence line defined by a thermodynamic first order phase transition Domb (2000). Nevertheless, the coexistence line has finite extension and ends into the so called critical point (e.g., $T_{c}\approx 647$ K and $P_{c}\approx 220$ bar for water), above which lies a large part of the phase diagram – the supercritical fluid (SCF) region Kiran et al. (2012); Proctor (2020) – which was first discovered by Baron de la Tour in 1822 de La Tour (1822). In this SCF region, liquid and gas phases can be connected without encountering any thermodynamic singularity, which raises the long-standing question of what really separates a liquid from a gas Brazhkin and Trachenko (2012), if anything, and whether a supercritical fluid constitutes a distinct phase of matter.
A commonly considered scenario is that the SCF region contains liquid-like and gas-like states separated in the phase diagram by a single separation line. Importantly, this line does not correspond to any thermodynamic singularity and therefore represents a continuous crossover, making both its definition and precise location inherently subtle. Indeed, multiple distinct and non-coincident definitions of this crossover line have been proposed in the literature, including the Widom line, Frenkel line, and Widom–Fisher line, among others Xu et al. (2005); Jones and Walker (1956); Luo et al. (2014); Banuti et al. (2017); Corradini et al. (2015); Gallo et al. (2014); Fisher and Widom (1969); Vega et al. (1995); Evans et al. (1993); Brazhkin et al. (2012a, 2013); Bolmatov et al. (2013, 2014); Fomin et al. (2018); Nishikawa and Tanaka (1995); Matsugami et al. (2014); Ploetz and Smith (2019); Woodcock (2013, 2016); Brazhkin et al. (2012b); Huang et al. (2023); Ouyang et al. (2024).
An alternative scenario, summarized in Fig. 1, posits that the liquid and gas phases are separated by a finite crossover region, bounded by two boundary lines Banuti (2015); Maxim et al. (2021); Wang et al. (2021); Li and Jin (2024), thereby rendering the SCF in this region distinguishable from both a liquid and a gas. A two-boundary picture has been explored, for instance, in the Widom delta scenario Ha et al. (2018); Yoon et al. (2019); Ha et al. (2020), where the crossover lines are associated with the percolation loci of gas-like and liquid-like clusters. In this framework, however, the intermediate SCF region is not treated as a distinct state of matter but rather as a gas–liquid coexistence regime. Moreover, the definition of such clusters is often ambiguous and typically relies on machine-learning-based classifications Ha et al. (2018), lacking a definitive interpretation of the associated physical signatures (see however Asharchuk et al. (2025)).
On the other hand, more recent studies have established the thermodynamic origin of these two supercritical boundary lines ($L^{\pm}$ lines)  Li and Jin (2024) and their universal scalings in liquid-gas phase transitions Li and Jin (2024), Ising models Li and Jin (2024), quantum phase transitions Wang et al. (2024); Lv et al. (2024), charged black holes Wang et al. (2025) and liquid-liquid phase transitions Wang et al. (2025).
If, however, the SCF constitutes an independent state of matter, as implied by the second scenario Li and Jin (2024), then a natural question arises: what defines this state, and what physical signatures mark the crossing of the two supercritical boundaries? Here, we address this question by showing that the SCF exhibits unique structural, transport, and dynamical properties, fundamentally different from those of both liquids and gases. These differences produce clear, albeit non-critical, physical signatures when crossing each of the two supercritical boundaries, consistently observed in both simulations and experiments.
Figure 1: Phase diagram of argon. The critical point (CP) is located at $(T_{c},P_{c})\approx(151\,\text{K},48.55\,\text{bar})$ Mccain and Ziegler (1967). The liquid-gas coexistence line (solid black line) is obtained from the NIST database NIS . The gray dotted line is the critical isochore, which represents supercritical continuation of the coexistence line. The two black dashed lines are $L^{\pm}$ lines, which are computed using the NIST data and the thermodynamic criterion proposed in Li and Jin (2024), and represent the boundaries between liquid, gas and SCF phases.
Phase diagram and supercritical boundaries. We consider Ar as a prototypical simple system and perform extensive MD simulations (see Methods). In Fig. 1a, we present the phase diagram of Ar. The critical point is located at $(T_{c},P_{c})\approx(151\,\text{K},48.55\,\text{bar})$ Mccain and Ziegler (1967). The liquid-gas coexistence line (the black line in Fig. 1a) continues into the supercritical regime (the gray dotted line), which typically shows a high symmetry revealed by maxima of density fluctuations under constant temperature conditions Xu et al. (2005); Nishikawa and Tanaka (1995); Ploetz and Smith (2019). In this study, we regard all supercritical lines defined in this spirit as the continuation of the coexistence line, including the critical isochore, Widom line Xu et al. (2005), Nishikawa line Nishikawa and Tanaka (1995), and symmetry line Ploetz and Smith (2019). These lines locate slightly differently in the phase diagram (see Li and Jin (2024)), but all around the middle of the SCF phase. Importantly, in our framework, and different from previous interpretation, those are not considered as the “boundary line” between liquid-like and gas-like states.
As shown in Fig. 1a, we propose that the gas, SCF and liquid phases are separated by two boundary lines, (black dashed lines). (i) The liquid-SCF boundary roughly corresponds to the Fisher-Widom line Vega et al. (1995), defined based on the behavior of the radial distribution function $g(r)$ (a structural viewpoint), or the Frenkel line defined based on dynamical viewpoints. The location of the Frenkel line depends on the specific criterion employed: the Frenkel-1 line is defined by the equality $\tau_{0}=\tau$ between the vibration time $\tau_{0}$ and the liquid relaxation time $\tau$ Brazhkin et al. (2012a), the Frenkel-2 line corresponds to the disappearance of oscillations and minima of the velocity autocorrelation function Brazhkin et al. (2013), and the Frenkel-3 line is derived based on isochoric heat capacity $C_{V}$ (for monatomic systems such as argon, the criterion is $C_{V}=2k_{\rm B}$Brazhkin et al. (2012a). (ii) The gas-SCF boundary, which is the focus of the present study, has never been individually investigated in previous works. (iii) Several existing considerations have proposed two boundaries, including the $L^{\pm}$ lines Li and Jin (2024), percolation lines Woodcock (2013, 2016), the two boundaries of the pseudo-boiling transitional region Banuti (2015); Wang et al. (2021), and the Widom delta Ha et al. (2018); Yoon et al. (2019); Ha et al. (2020). Different criteria give nearby, but quantitatively different boundary lines (see Li and Jin (2024)). Again, the reason is that these lines do not correspond to any thermodynamic or dynamic singularities, which means that they cannot be uniquely defined in a mathematically rigorous way even in the thermodynamic limit. An exception is the $L^{\pm}$ lines in quantum phase transitions, where the thermodynamic supercritical crossovers coincide with dynamic singularities under quantum quenches Wang et al. (2024).
Importantly, among the above proposals, only the $L^{\pm}$ lines display universal scalings in various phase transitions Li and Jin (2024); Wang et al. (2024); Lv et al. (2024); Wang et al. (2025): the scalings of the external field and order parameter along $L^{\pm}$ lines satisfy $\delta P^{\pm}\sim(T-T_{\rm c})^{\beta+\gamma}$ and $\delta\rho^{\pm}\sim(T-T_{\rm c})^{\beta}$, where $\beta$ and $\gamma$ are standard critical exponents. Because of these reasons, below, we adopt the $L^{\pm}$ lines, introduced in Ref. Li and Jin (2024) and shown for the case of argon as black dashed curves in Fig. 1, as the reference for the two supercritical boundaries (see Methods).
For the detailed analyses below, we select two isothermal lines corresponding to $T=195$ K (results in the main text) and $T=220$ K (see Supplementary Material Sec. S1) respectively and consider a large range of pressures that scan from largely above the liquid-SCF boundary to well below the gas-SCF boundary. This will allow us to study in detail the structural, transport and dynamical properties across the boundaries, revealing their fundamental origin and nature.
Figure 2: Radial distribution functions (RDFs). (a-c) Typical reduced RDF, $\tilde{h}(r)=r[g(r)-1]$, of gas, SCF and liquid states measured in simulations of argon (gas: $T=195$ K, $P=20~\text{bar}$; SCF: $T=195$ K, $90~\text{bar}$; liquid: $T=100$ K, $P=40~\text{bar}$). The distance $r$ is expressed in units of $\sigma_{0}$, the position of the first peak. (d) Radial distribution function $g(r)$ at different pressures, zoomed in around the second peak, for T=195KT=195\ $\text${K}. (inset) Height of the second peak $h_{2}$ as a function of $P$; below $P_{2}\approx 55$ bar (red circle), the second peak can not be identified anymore.
Structural evidence: radial distribution functions. In Fig. 2(a-c), we present the reduced radial distribution function, $\tilde{h}(r)=r[g(r)-1]$, for argon in three distinct states: gas, SCF, and liquid. The gas state exhibits a rapid decay of the tail of $\tilde{h}(r)$ without oscillations, indicating the weak intermolecular correlations. In contrast, the liquid displays exponentially damped (yet persistent) oscillations with multiple distinct peaks, signaling robust short-range order. The SCF exhibits an intermediate structure: a few peaks emerge, yet higher-order coordination shells are absent, indicating loss of orders at large distances. This represents a unique structural signature of the SCF, setting it apart from the gas and liquid states.
Based on above evidence, we propose that the structure of SCF is characterized by sub-short-range (SSR) order. The SSR order is reflected in the radial distribution function $g(r)$ through damped oscillations that are truncated, i.e., the oscillatory behavior of $g(r)$ extends only up to a finite SSR length scale. This viewpoint on structure and, as we will see, also the behavior of transport and dynamical properties, imply that the SCF can be interpreted as an independent state of matter in a loose sense. This definition is not rigorous because there are no bona fide phase transitions between SCF and gases/liquids and all lines should be rather regarded as smooth crossovers. Nevertheless, these three phases can be still separated by two supercritical boundary lines:
(i) Liquid-SCF boundary. From a structural perspective, the Fisher-Widom line Fisher and Widom (1969) corresponds to the liquid-SCF boundary. In liquids, the oscillations in $g(r)$, $\sim\cos(qr)$, although damped, remain up to $r\to\infty$. The liquid-SCF boundary has been alternatively estimated based on dynamic criteria (Frenkel line Brazhkin et al. (2012a)), such as the disappearance of the short-time vibrational motion typical of liquids. The Fisher-Widom line and the Frenkel line are also close to the upper $L^{+}$ line defined from thermodynamic and scaling considerations Li and Jin (2024). In general, the liquid-SCF boundary corresponds to the boundary beyond which the state loses liquid characteristics, although its precise location depends on the detailed criteria. The small quantitative inconsistency is possibly inevitable (also for the gas-SCF boundary), due to the absence of thermodynamic or dynamic singularities. Nevertheless, these lines are close to each other, and can be unambiguously distinguished from the gas-SCF boundary (as we will show in this study). Since the liquid–SCF transition has been extensively studied in the literature, including from the perspective of structural criteria such as the Fisher-Widom line Fisher and Widom (1969); De Carvalho et al. (1994); Vega et al. (1995); Dijkstra and Evans (2000); Evans et al. (1993); Tarazona et al. (2003); Stopper et al. (2019), we do not focus on it here. Instead, our study concentrates on the gas-SCF boundary, which remains largely unexplored.
(ii) Gas-SCF boundary. This boundary corresponds to the locus where the oscillatory behavior in $g(r)$, aside from the first peak which persists except in the ideal gas limit, disappears. In practice, it can be located by the disappearance of the second peak in $g(r)$, or equivalently the first minimum, as shown by the transition from Fig.1(b) to Fig.1(c), using MD simulation data of argon.
Figure 3: Transport coefficients. (a) Reduced viscosity $\eta^{*}=\eta\left(\sigma^{2}/\sqrt{\epsilon m}\right)$, where $\sigma$ and $\epsilon$ are length and energy parameters in the Lennard-Jones potential (see Methods), and $m$ is the atomic mass. The errors bars represent the standard error of the mean. (b) Reduced diffusion coefficient $D^{*}=D\sqrt{\left(\frac{m}{\epsilon\sigma^{2}}\right)}$. In the kinetic theory expression, $D_{0}=\frac{3}{8\rho\sigma^{2}}\left(\frac{k_{\rm B}T}{\pi m}\right)^{1/2}$, we have used both the ideal gas form $\rho_{\rm ideal}(P,T)=k_{\rm B}P/T$ (ideal kinetic theory), and the exact $\rho(P,T)$ from simulation data (kinetic theory). (c) Reduced diffusion coefficient as a function of the two-particle structural entropy $S_{2}$. The blue dashed line represents an exponential fit of the data above the $L^{+}$ line, $D=1.18e^{S_{2}/k_{\rm B}}$.
More specifically, we measure $g(r)$ using MD simulations at $T=195$ K, in the pressure range from 5 bar to 900 bar (see Fig. S1a) and we focus on the second peak in $g(r)$ (see Fig. 2d). Based on the data of $g(r)$, a separation pressure $P_{2}\approx 55$ bar is determined: when $P>P_{2}$, the height of the second peak $h_{2}(P)$ increases with $P$; when $P<P_{2}$, the second peak cannot be identified within the given accuracy of the numerical data.
Below $P_{2}$, the $g(r)$ has only one peak, $g(r)\approx\exp[-V(r)/k_{\rm B}T]$, where $V(r)$ is the pair interparticle interaction, as given by the low-density theory Chandler (1987). In this case, $g(r)$ decays beyond the first peak without any oscillations – this is the definition of the gas state. On the other hand, $P_{2}$ should correspond to the gas-SCF boundary at $T=195$ K. Interestingly, $(P_{2}\approx 55\,{\rm bar}$, $T=195\,~\rm{K})$ coincides with the point $(P^{-}\approx 60\,{\rm bar}$, $T=195\,\rm{K})$ on the $L^{-}$ line estimated independently in  Li and Jin (2024) using thermodynamics, confirming $L^{-}$ as the gas-SCF boundary. In the Supplementary Material, we verify the universality of these findings by considering a second isothermal cut crossing the $L^{\pm}$ lines (see Fig. S2). This observation suggests that, from a structural perspective, the $L^{-}$ line corresponds to the disappearance of the second peak in $g(r)$.
Below, we reveal that the supercritical boundaries are also reflected in transport properties, specifically, the viscosity $\eta$ and the self-diffusion coefficient $D$, as well as in the dynamical behavior, such as the velocity autocorrelation function (VACF), $Z(t)$. These observations can also be derived from a liquid-state theory based on the memory-function approach and are supported by simulation results. The numerically estimated gas–SCF boundary based on these criteria aligns closely with the lower $L^{-}$ line independently identified in Ref. Li and Jin (2024) using a thermodynamic criterion, suggesting a clear physical significance for the latter.
Kinetic features: transport coefficients. Next, we present clear signatures of the $L^{\pm}$ crossover lines in the transport coefficients. In the dilute limit, transport processes are well described by the well-known kinetic theory Chapman and Cowling (1990); Hansen and McDonald (2013). However, beyond the gas-SCF boundary, non-negligible corrections to the transport coefficients emerge - in particular, to the viscosity $\eta$, where the excess part $\eta_{\rm ex}$ that depends on the interaction potential and configurational structure becomes dominant compared to the kinetic part $\eta_{0}$.
Figure 4: Velocity auto-correlation functions. (a) Fitting the tail of the VACF. Assuming a stretched exponential form, $Z(t)\sim e^{-(t/\tau)^{\beta}}$, the exponent $\beta$ is obtained by the slope of the linear fit in the log-log plot of $\ln[Z(t)]$ vs. $t$. (b) $\beta$ as a function of $P$. Error bars represent the standard deviation of the fitted values from ten independent fitting ranges. (c) Einstein frequency $\Omega$ (Thz) as a function of pressure $P$. The power-law fits (dashed colored lines) highlight the two distinct behavior in the gas and liquid states.
We focus on the diffusion coefficient $D$ and viscosity $\eta$. In MD simulations, $D$ and $\eta$ are computed respectively using the VACF and the stress autocorrelation function, based on the Green-Kubo formula (see Methods for details). For comparison, we also collect the viscosity data from the National Institute of Standards and Technology (NIST) database NIS , which are consistent with our simulation data within numerical accuracy.
To analyze the simulation results and NIST data (see Fig. 3), we first examine the kinetic theory expressions, $D_{0}=\frac{3}{8\rho\sigma_{0}^{2}}\left(\frac{k_{\rm B}T}{\pi m}\right)^{1/2}$ and $\eta_{0}=\frac{5}{16\sigma_{0}^{2}}\left(\frac{mk_{\rm B}T}{\pi}\right)^{1/2}$, which are valid in the dilute limit Chapman and Cowling (1990); Hansen and McDonald (2013). As shown in Fig. 3a, kinetic theory describes well the viscosity data below the $L^{-}$ line, where $\eta$ is a constant independent of pressure, as expected for gases. However, kinetic theory significantly fails above that line, as $\eta$ exhibits a rapid increase with pressure. The location where this happens is in very good agreement with the thermodynamic estimate of the $L^{-}$ crossover line Li and Jin (2024).
To understand the origin of this deviation, one should note the approximations employed within such a formalism. Kinetic theory, as indicated by its name, only considers the kinetic effects, ignoring any explicit contributions from the interparticle potential. The theory also neglects the memory effects in particle motions by assuming that successive collisions are completely uncorrelated. Mathematically, this means that the memory function of the stress autocorrelation function $G(t)$ is given by $M_{G}(t)\sim\delta(t)$, where $\delta(t)$ is the delta function (see Supplementary Material Sec. S2). As we explicitly demonstrate, both approximations fail above the $L^{-}$ line. Moreover, it is important to emphasize that the microscopic mechanisms underlying fluid viscosity differ fundamentally between gases and liquids. In gases, viscosity arises from particle collisions, whereas in liquids it is governed by potential energy barriers, processes that kinetic theory does not account for. This supports our proposal that the $L^{-}$ line marks the crossover where potential effects become significant and the dynamics are no longer fully described by pair collisions, as in the gaseous regime.
In general, the viscosity contains two terms, $\eta=\eta_{0}+\eta_{\rm ex}$, where $\eta_{0}$ is the kinetic contribution predicted by the kinetic theory, and $\eta_{\rm ex}$ comes purely from the interaction potential. Using this separation, the gas-SCF crossover line can be formally identified as the location at which $\eta_{\rm ex}\approx 0$ no longer holds. Lucks and Moser developed a formalism (hereafter referred to as LM theory) to compute $\eta_{\rm ex}$ based on a memory-function approach Lucas and Moser (1979a, b). In order to achieve a more quantitative description, we take the version of the LM theory that approximates the excess memory function by a simple exponential form $M_{G}(t)=ae^{-bt}$, with $a$ and $b$ explicitly related to $g(r)$ and $V(r)$ (see Supplementary Material Sec. S2). As $P$ increases above $L^{-}$, $\eta=\eta_{0}+\eta_{\rm ex}$ quickly exceeds $\eta_{0}$. Using $g(r)$ from simulations, the theoretical values of $\eta$ computed using the LM theory are very close to simulation and NIST data (see Fig. 3a). Surprisingly, LM theory works reasonably well for both $\eta$ and $D$ even above the $L^{+}$ crossover lines. It remains unclear whether this is merely a fortuitous coincidence.
Among transport properties, the diffusion coefficient $D$ presents a more complex case since it is influenced by the structural order in a more indirect way. Because $D$ is related to the integral of the VACF, it does not have an excess term, in addition to the kinetic expression, $D_{0}=\frac{3}{8\rho\sigma_{0}^{2}}\left(\frac{k_{\rm B}T}{\pi m}\right)^{1/2}$. However, because this expression contains the equation of state (EOS) $\rho(P,T)$, which is connected to $g(r)$ and $V(r)$ through the virial equation, the kinetic expression $D_{0}$ depends on structure as well, despite in a more subtle way.
Indeed, $D_{0}$ computed using the full EOS agrees well with simulation data even beyond the $L^{-}$ line, extending approximately up to the $L^{+}$ line. In contrast, if the ideal gas equation of state is used, i.e., $\rho_{\rm ideal}(P,T)=P/(k_{\rm B}T)$, the resulting “ideal kinetic theory” begins to fail around the $L^{-}$ line (see Fig. 3b). This further supports our interpretation of $L^{-}$ as the boundary where structural correlations beyond nearest neighbors become significant and the system departs from gas-like behavior. Moreover, this analysis highlights that the gas–SCF boundary can be identified through transport coefficients, whose behavior is influenced by the emergence of SSR structural order in the SCF phase.
As already emphasized, above the liquid-SCF boundary (Frenkel-line), transport properties are dominated by a different microscopic mechanism. Meta-stable states in the potential landscape are formed, and the transitions between such states are realized by activation processes. In this regime, the Arrhenius law gives a relation between $D$ and the typical activation energy barrier $\Delta E$, $D\sim e^{-\Delta E/k_{\rm B}T}\sim e^{{\Delta S}/k_{\rm B}}$. The excess entropy $\Delta S$} can be approximately described by the two-particle entropy $S_{2}=-2\pi\rho k_{\rm B}\int_{0}^{\infty}\{g(r)\ln[g(r)]-[g(r)-1]\}r^{2}dr$ Dzugutov (1996), which can be directly computed from the $g(r)$ data. Leveraging on these arguments, we find that the scaling $D\sim e^{{S_{2}}/k_{\rm B}}$ deviates from the simulation data around $L^{+}$, as expected (see Fig. 3c) Demmel and Jakse (2025).
Dynamical fingerprints: velocity auto-correlation function (VACF). Besides the signatures of the SCF state in structural and transport properties, the supercritical boundaries can be also identified by pure dynamic probes. We focus on the normalized VACF $Z(t)$ (Fig. S1b), which satisfies, $\dot{Z}(t)=-\int_{0}^{t}M_{Z}(t-\tau)Z(\tau)d\tau$, where $M_{z}(t)$ is the memory function. Interestingly, the intermediate-time behavior of the VACF has been used to locate the liquid-SCF boundary (the Frenkel-2 line). In fact, at high pressures, the VACF displays a minimum after a fast decay; this minimum is gradually suppressed by decreasing pressure, which eventually disappears around the crossover pressure defined by the Frenkel-2 line Brazhkin et al. (2013).
Here we propose that the large-time behavior of the VACF, but also the very short one incarnated in the so-called Einstein frequency, contain information that reflects the gas-SCF boundary. It can be shown that (see Supplementary Material Sec. S2), when the memory function is $M_{Z}\sim\delta(t)$ (memory-less) or exponential $M_{Z}(t)\sim e^{-t/\tau_{M}}$, the VACF decays exponentially at large times, $Z(t)\sim e^{-t/\tau}$. However, this simple exponential form is consistent with our simulation data only below the $L^{-}$ line. Above the $L^{-}$ line, we find that at late time (up to the precision of our numerical data) the VACF is well-approximated by a stretched exponential, $Z(t)=A\exp\left[\left(-\frac{t}{\tau}\right)^{\beta}\right]$, with $\beta<1$. In Fig. 4b we show the exponent $\beta$, obtained by fitting the data with the stretched exponential form (Fig. 4a), as a function of pressure (see Supplementary Material Sec. S3 for additional details). We find that the $L^{-}$ line coincides with the location at which $\beta$ approaches one. The stretched exponential of VACF (with $\beta<1$) suggests the onset of non-trivial memory effects beyond $L^{-}$, but its theoretical derivation deserves to be investigated in the future.
In addition, the signature of $L^{\pm}$ lines can also be found in the Einstein frequency $\Omega$ that appears in the very-short-time expansion of $Z(t)=1-\Omega^{2}\frac{t^{2}}{2}+\cdots$ Hansen and McDonald (2013) (see Fig. 4c). These results imply that, beyond structural and transport properties, the dynamical behavior of the SCF state also clearly distinguishes it from both the gas and liquid phases, reinforcing its identification as an intermediate state of matter.
Figure 5: Experimental diffusion coefficient for supercritical methane at $T=200$ K. Experimental data obtained by quasi-elastic neutron scattering (QENS) Ranieri et al. (2024) and nuclear magnetic resonance (NMR) Oosting and Trappeniers (1971) are shown respectively with open circles and open diamonds. The green line is the theoretical prediction from ideal kinetic theory, with scaling $D^{*}\propto P^{-1}$.
Experimental signatures and proposals. Our findings provide a basis for experimental verification in future studies. Accurate measurements of $g(r)$ in the SCF state are becoming available using modern neutron scattering instruments Dove and Li (2022); Iwase et al. (2010) and the state-of-the-art data analysis tools Ma et al. (2020).
It is also possible to measure transport coefficients using experimental techniques such as neutron scattering. In fact, a recent experimental study has already reported accurate data of $D$ in supercritical methane along the $T=200$ K isotherm Ranieri et al. (2024). In Fig. 5, we present the experimental data from Refs. Oosting and Trappeniers (1971); Ranieri et al. (2024), together with the $L^{\pm}$ lines determined based on the NIST EOS of methane. The experimental data are fully consistent with our theoretical expectation. The diffusion constant $D$ follows the ideal kinetic theory behavior $D\sim 1/P$ up to a pressure around the $L^{-}$ line as in Fig. 3b. Beyond the $L^{-}$ line a clear deviation from this scaling is observed. The SCF state between $L^{\pm}$ lines appears to coincide with a “transition regime” between the gaseous low-$P$ and the liquid high-$P$ behavior. In agreement with our simulation results, this observation provides a first experimental evidence of the two $L^{\pm}$ crossover lines and of the existence of the supercritical fluid state, clearly separated from the gas and liquid phases. During the finalization of our work, Ref. Khrapak et al. (2025) appeared, presenting a parallel theoretical analysis of the experimental data from Ref. Ranieri et al. (2024). At present, any connection between our results and the ideas put forward in Ref. Khrapak et al. (2025) remains unclear, but certainly warrants further investigation.
Finally, the signatures of the supercritical boundary lines could be experimentally observed in a more phenomenological way. Enhanced density fluctuations have been observed in small-angle neutron scattering experiments around the $L^{+}$ line Li and Jin (2024) in supercritical carbon dioxide Pipich and Schwahn (2018), due to the formation of liquid-like droplets. Analogously, we also expect enhanced density fluctuations around the $L^{-}$ line, due to formation of gas-like bubbles. If one interprets the supercritical transition regime between $L^{\pm}$ lines as the blur-up of the subcritical discontinuous first-order transition, then the rapid change of the EOS in this regime might be related to the phenomenon of supercritical pseudo-boiling Banuti (2015); Wang et al. (2021); He et al. (2023).
Discussion. In this study, we demonstrate that the signatures of the supercritical fluid boundaries can be identified by various physical quantities, including structure descriptors such as the radial distribution function, transport coefficients, and the velocity auto-correlation function. Our study completes the phase diagram above the critical point clarifying the existence of three independent states of matter characterized by distinct physical properties. (i) The liquid-SCF boundary is the limit beyond which the system does not behave like a typical liquid. Because the liquid state can be characterized by various features, ranging from thermodynamics to dynamics to structures, the deviation from the liquid behavior can be defined according to different criteria Fisher and Widom (1969); Brazhkin et al. (2012a, 2013, a). (ii) The gas-SCF boundary is the limit beyond which the system does not behave like a typical gas. This line is reflected in several physical properties and can be therefore defined using thermodynamics, structural order, transport coefficients and finally dynamical behavior. (iii) In-between is the SCF state that is neither liquid-like nor gas-like but rather a separate state of matter, whose properties cannot be easily rationalized as a simple linear superposition of gas and liquid features. We find that, the Widom line defined by the maximum of the Ornstein-Zernike correlation length (Supplementary Material Sec. S4Xu et al. (2005), and the Widom delta identified according to the fractions of liquid-like and gas-like atoms (Supplementary Material Sec. S5Jun et al. (2018), both lie within the intermediate region bounded by $L^{\pm}$ lines.
It remains unclear whether our findings, and in particular the existence of a gas-SCF boundary, extend to systems with purely repulsive interactions, where no thermodynamic separation between liquid and gas phases, and no critical point, exist. So far, in these systems, only a boundary line between gas-like and liquid-like states has been discussed, e.g., Huang et al. (2023); Jiang et al. (2025). Moreover, whether our results extend to more complex molecular fluids such as CO2 or H2O remains an open question and warrants further investigation.
On the other hand, the proposed framework and analysis methods can be generalized to many other systems. In particular, it would be very interesting to study the nature of order in the quantum supercritical state Wang et al. (2024); Lv et al. (2024), in the context of quantum many-body correlations. Moreover, the dynamical signatures of the $L^{-}$ line can be explored in the supercritical kinetics of black holes Zhao et al. (2025a); Li et al. (2025a, b) and holographic models as well Zhao et al. (2025b). Another future direction is to examine in more detail the analogy between the supercritical fluid state and the QCD phase diagram of nuclear matter. Interestingly, a three-state phase diagram presenting an intermediate partially deconfined phase, and naive similarities with our proposal depicted in Fig. 1(a), has been recently discussed Hanada and Watanabe (2023); Glozman et al. (2022); Fujimoto et al. (2025).
In conclusion, returning to the fundamental question of “what separates a liquid from a gas” Brazhkin and Trachenko (2012), our revised answer points to the existence of an intermediate supercritical fluid state, clearly distinguishable from both gas and liquid phases in terms of thermodynamic, structural, transport, and dynamical properties.
Methods
Molecular dynamics simulations
In MD simulations, we employ the Lennard-Jones (LJ) potential for argon in the LAMMPS package Thompson et al. (2022), which benefits from efficient parallelization for large atom counts. The expression for the LJ potential is given as
$V(r)=4\epsilon\left[\left(\frac{\sigma}{r}\right)^{12}-\left(\frac{\sigma}{r}\right)^{6}\right],$ (1)
where $\epsilon/k_{\rm B}=119.5$ K, $\sigma=$3.405\text{\,}\mathrm{\text{\AA }}$\$, and $k_{\rm B}=$1.38\text{\times}{10}^{-23}\text{\,}\mathrm{J}\text{\,}{\mathrm{K}}^{-1}$\$ is the Boltzmann constant. To address potential finite-size effects, we have evaluated systems of 500, 4,000, and 32,000 atoms. Our analysis confirm that 4,000 atoms provides a balance between statistical accuracy and computational efficiency for calculating transport properties; consequently, this system size is adopted for the production runs. The initial system is a face-centered cubic lattice of 4000 atoms with a lattice constant of $5.26\text{\,}\mathrm{\text{\AA }}$. Thirty-five different pressures are simulated from $5\text{\,}\mathrm{bar}$ to $1300\text{\,}\mathrm{bar}$. For each target pressure, the system is initially relaxed at $195\text{\,}\mathrm{K}$ for $5\text{\,}\mathrm{ns}$ in an isothermal-isobaric (NPT) ensemble to reach thermal equilibrium. Following the NPT stage, the system undergoes an intermediate 500 ps relaxation in the microcanonical (NVE) ensemble at each specific state point to ensure total energy stability and facilitate unconstrained dynamical relaxation. Finally, a canonical (NVT) relaxation is performed to stabilize the system prior to the $20\text{\,}\mathrm{ns}$ production phase – all measurements are performed in this time period. The temperature and pressure are regulated via the Nosé-Hoover methods Nosé (1984); Hoover (1985), with Newton’s equations of motion solved numerically using a time step of $1\text{\,}\mathrm{fs}$. All samples are simulated with periodic boundary conditions. The pressure unit is bar, length unit Angstroms, time unit fs, and energy unit kcal/mol.
The diffusion coefficient $D$ is calculated using the Einstein relation from the mean square displacement (MSD) of a 15 ns NVT equilibrium trajectory at each pressure. The viscosity $\eta$ is calculated using the Green–Kubo formalism, which relates $\eta$ to the integral of the stress autocorrelation function. Specifically, $\eta$ is given by
$\begin{split}\eta=\frac{1}{Vk_{\rm B}T}\int_{0}^{\infty}\left\langle S^{\alpha\beta}(0)S^{\alpha\beta}(t)\right\rangle dt,\\ S^{\alpha\beta}=-\sum_{i}^{N}\left(mv_{i}^{\alpha}v_{i}^{\beta}+f_{i}^{\alpha}r_{i}^{\beta}\right),\end{split}$ (2)
where $S^{\alpha\beta}(t)$ ($\alpha\neq\beta$) denotes an off-diagonal component of the microscopic stress tensor, expressed as the sum of kinetic ($mv_{i}^{\alpha}v_{i}^{\beta}$) and virial ($f_{i}^{\alpha}r_{i}^{\beta}$) contributions. In our simulations, the stress autocorrelation function is computed from a long-time NVT trajectory of at least 15 ns. The time correlation is averaged over multiple time origins along the trajectory, ensuring the convergence of the Green–Kubo integral within the correlation time. For low-pressure systems, longer trajectories are used to capture the slower decay of stress fluctuations, and the average over multiple independent simulations with different initial conditions are performed to further reduce the statistical noise.
Thermodynamic definition of $L^{\pm}$ lines
To define from thermodynamics the $L^{\pm}$ lines Li and Jin (2024), one first chooses the critical isochore, which is an extension of the coexistence line to the supercritical region, as the reference line. The compressibility $\kappa_{T}\equiv\frac{P_{c}}{\rho^{\rm m}_{c}}\left(\frac{\partial\rho^{\rm m}}{\partial P}\right)_{T}$ is evaluated along each path parallel to the critical isochore, where $P_{c}$ and $\rho^{\rm m}_{c}$ are the critical pressure and critical mass density. Since $\kappa_{T}$ is a function of distance $\delta P(\rho^{\rm m},T)=P(\rho^{\rm m},T)-P(\rho^{\rm m}_{c},T)$ and $T$, one can find a temperature $T_{\rm max}(\delta P)$ that maximizes $\kappa_{T}$ along each path with a fixed $\delta P$. All the $T_{\rm max}(\delta P)$ points under different $\delta P$ together consist of $L^{\pm}$ lines. In order to determine $L^{\pm}$ lines, one needs to know the EOS $P(\rho^{\rm m},T)$. In this study, we take the EOS of argon from the NIST database NIS . The resulting $L^{\pm}$ lines are plotted in Fig. 1a.
Acknowledgements.
We thank S.Khrapak for bringing the recent Ref. Khrapak et al. (2025) to our attention. SJ and MB acknowledge the support of the Shanghai Municipal Science and Technology Major Project (Grant No.2019SHZDZX01). XL acknowledges the Postdoctoral Fellowship Program of CPSF (Grant Number GZC20252776). XF acknowledges the support of the National Natural Science Foundation of China (No.52201016). MB acknowledges the support of the sponsorship from the Yangyang Development Fund. YJ acknowledges funding from the National Key R&D Program of China (Grant No. 2025YFF0512000), Project supported by the Space Application System of China Manned Space Program, NSFC (Grant NO. 12447101) and Wenzhou Institute (Grant WIUCASICTP2022). SJ and XF thank the Key State Laboratory for Clean and Efficient Turbomachinery Power Equipment for providing High Performance Cluster resources. The authors acknowledge the use of the High Performance Cluster at Institute of Theoretical Physics, Chinese Academy of Sciences, and the computer clusters at the Hefei advanced computing center.
Supplementary Information
In the supplementary information we provide extended data, details of the analyses presented in the main text and additional technical information about the theoretical frameworks used.
S1 Supplementary analysis of radial distribution functions (RDFs).
In Fig. S1, we plot a complete set of RDFs measured by molecular dynamics (MD) simulations at $T=195$ K and different pressures. The radial distribution functions are calculated by averaging over a large number of frames to reduce statistical noise. We extract 500 ps of trajectories with 1 fs intervals, resulting in 500,000 frames for averaging. These extensive averages ensure that the RDFs are well-converged.
In Fig. S2, we show RDFs enlarged around the second peak, at another temperature, $T=222$ K. The second peak emerges around $P_{2}=55$ bar, consistent with the pressure $P^{-}=47$ bar at the $L^{-}$ line. This test provides an additional verification of the structural signature at the $L^{-}$ line.
Next, we discuss the procedure employed to determine the peaks in $g(r)$. To quantitatively track the evolution of structural features in the radial distribution function $g(r)$, its second peak is fitted using a Gaussian function. Specifically, we employ a Gaussian function of the following form to fit the second peak,
$f(r)=c+h_{2}\exp\!\left[-\frac{(r-r_{0})^{2}}{2\sigma^{2}}\right],$ (S1)
with fitting parameters $c,h_{2},r_{0}$ and $\sigma$. Here $h_{2}$ and $r_{0}$ provide the height of position respectively. When the coefficient $R^{2}$ of determination is below $0.8$, the Gaussian model no longer provides an adequate description of the feature, indicating that the second peak cannot be reliably identified.
Figure S1: Radial distribution functions at T=195KT=195\ $\text${K}. Radial distribution functions $g(r)$ for argon are obtained from MD simulations at P=5,15,25,35,45,55,60,75,90,105,120,135,150,200,320,500,700,800barP=5,15,25,35,45,55,60,75,90,105,120,135,150,200,320,500,700,800\ $\text${bar} (from red to blue) and T=195KT=195\ $\text${K}.
Figure S2: Enlarged radial distribution functions at T=220KT=220\ $\text${K}. Radial distribution function $g(r)$ at different pressures, zoomed in around the second peak, for T=220KT=220\ $\text${K}.
S2 Theories of auto-correlation functions and transport coefficients
S2.1 Velocity auto-correlation function (VACF) and diffusion coefficient
The (normalized) VACF is defined by,
$Z(t)\equiv\frac{\left\langle{\bf v}(0)\cdot{\bf v}(t)\right\rangle}{\left\langle{\bf v}(0)\cdot{\bf v}(0)\right\rangle},$ (S2)
and the diffusion coefficient $D$ is related to the VACF through the Green-Kubo formula,
$D=\frac{k_{\rm B}T}{m}\int_{0}^{\infty}Z(t)dt,$ (S3)
where $T$ is the temperature, $m$ the mass of particle, $k_{\rm B}$ the Boltzmann constant. In general, the VACF satisfies the following equation,
$\dot{Z}(t)=-\int_{0}^{t}M_{Z}(t-\tau)Z(\tau)d\tau,$ (S4)
where $M_{Z}(t)$ is the memory function describing the memory effects in the motion of particles, and $\dot{Z}$ represents the time derivative of $Z$.
S2.1.1 Kinetic theory
In the dilute limit, the dynamics are memory-less,
$M_{Z}(t)=\frac{1}{\tau}\delta(t),$ (S5)
and the VACF decays exponentially,
$Z(t)=e^{-t/\tau},$ (S6)
where $\tau$ is the velocity correlation time. Equation (S3) then gives,
$D=\frac{k_{\rm B}T\tau}{m}.$ (S7)
Elementary kinetic theory Chapman and Cowling (1990) shows that in the dilute limit, $\tau=l/\langle v\rangle$, where $l$ is the mean free path and $\langle v\rangle$ the average velocity. For hard spheres, the kinetic theory gives, $\tau=\frac{3}{8\rho\sigma_{0}^{2}}\left(\frac{\pi k_{\rm B}T}{m}\right)^{-1/2}$, where $\sigma_{0}$ is the diameter of hard spheres, and
$D_{0}=\frac{3}{8\rho\sigma_{0}^{2}}\left(\frac{k_{\rm B}T}{\pi m}\right)^{1/2}.$ (S8)
For the Lennard-Jones potential, $r=\sigma_{0}\approx 1.12\sigma$ is where the potential energy $V(r)$ (see Eq. 1) minimizes.
S2.1.2 Lucks-Moser theory
At high densities, the memory-less assumption, Eq. (S5), does not hold anymore. However, in the interested regime, i.e., for the SCF around the gas-SCF boundary, the memory effects are not significant, and the memory function $M_{Z}(t)$ decays much faster than the VACF $Z(t)$. In this case, one can assume a simple form of $M_{Z}(t)$, such as the following exponential form,
$M_{Z}=\Omega^{2}e^{-t/\tau_{M}}.$ (S9)
Here $\Omega$ is the Einstein frequency that appears in the short-time expansion of $Z(t)=1-\Omega^{2}\frac{t^{2}}{2}+\cdots$, and $\tau_{M}$ is the characteristic time of the memory function. Plugging Eq. (S9) into into Eq. (S4), one obtains,
$Z(t)=\frac{1}{\alpha_{+}-\alpha_{-}}\left(\alpha_{+}e^{-\alpha_{-}t}-\alpha_{-}e^{-\alpha_{+}t}\right),$ (S10)
where
$\alpha_{\pm}=\frac{1}{2\tau_{M}}\left[1\mp\left(1-4\Omega^{2}\tau_{M}^{2}\right)^{1/2}\right].$ (S11)
Equation (S3) then gives,
$D=\frac{k_{\rm B}T}{m\Omega^{2}\tau_{M}}.$ (S12)
In order to explicitly compute $D$, one needs to know $\Omega^{2}$ and $\tau_{M}$. Lucks and Moser has related these two quantities to $g(r)$ and the inter-particle potential $V(r)$, through the following formulas Lucas and Moser (1979a):
$\Omega^{2}=\frac{\left\langle\dot{\bf v}(0)\cdot\dot{\bf v}(0)\right\rangle}{\left\langle{\bf v}(0)\cdot{\bf v}(0)\right\rangle},$ (S13)
$\frac{1}{\tau_{M}}=\left[\frac{\left\langle\ddot{\bf v}(0)\cdot\ddot{\bf v}(0)\right\rangle}{\left\langle\dot{\bf v}(0)\cdot\dot{\bf v}(0)\right\rangle}-\frac{\left\langle\dot{\bf v}(0)\cdot\dot{\bf v}(0)\right\rangle}{\left\langle{\bf v}(0)\cdot{\bf v}(0)\right\rangle}\right]^{1/2},$ (S14)
where
$\left\langle{\bf v}(0)\cdot{\bf v}(0)\right\rangle=\frac{3k_{\rm B}T}{m},$ (S15)
$\frac{\left\langle\dot{\bf v}(0)\cdot\dot{\bf v}(0)\right\rangle}{3k_{\rm B}T/m\sigma_{0}}=32\pi\frac{n^{*}}{T^{*}}\int_{0}^{\infty}g_{2}\left(r_{12}^{*}\right)\left[3W_{12}^{\prime}+2W_{12}^{\prime\prime}r_{12}^{*2}\right]r_{12}^{*2}dr_{12}^{*},$ (S16)
[…]
with
$\begin{aligned} W_{ij}=\frac{V(r)}{4\epsilon},\quad W_{ij}^{\prime}=dW_{ij}/dr_{ij}{}^{*2},\quad W_{ij}^{\prime\prime}=d^{2}W_{ij}/\left(dr_{ij}{}^{*2}\right)^{2},\quad n^{*}=n\sigma_{0}^{3},\quad T^{*}=\frac{k_{\rm B}T}{\epsilon}.\end{aligned}$ (S18)
Equations (S12)-(LABEL:eq:ddv2) give a theoretical estimate of the diffusion coefficient, $D_{\rm LM}$. Compared to the kinetic theory $D_{0}$, $D_{\rm LM}$ takes into account the dynamical memory effects (Eq. S9) and the spatial pair correlations $g(r)$.
Unsurprisingly, even the Lucks-Moser theory breaks down if the density is further increased. The VACF Eq. (S10) still has an exponential form at large times, with a correlation time $\tau\sim 1/\alpha_{-}$ (note that $\alpha_{+}<\alpha_{-}$ ). As shown in the main text, this exponential form does not hold anymore at high densities – instead, we need a stretched exponential function to fit the simulation data. However, the exponential approximation does not seem to affect significantly the transport coefficients.
S2.2 Stress auto-correlation function and Viscosity
The viscosity is connected to the stress auto-correlation function,
$G(t)\equiv\left\langle S^{\alpha\beta}(0)S^{\alpha\beta}(t)\right\rangle,$ (S19)
via the Green-Kubo relation,
$\eta=\frac{1}{Vk_{\rm B}T}\int_{0}^{\infty}G(t)dt.$ (S20)
Here $S^{\alpha\beta}$ represents the $\alpha\beta$ component of the off-diagonal ($\alpha\neq\beta$) stress tensor, defined as,
$S^{\alpha\beta}=-\sum_{i}^{N}\left(mv_{i}^{\alpha}v_{i}^{\beta}+f_{i}^{\alpha}r_{i}^{\beta}\right)=S^{\alpha\beta}_{0}+S_{\rm ex}^{\alpha\beta},$ (S21)
where $f_{i}^{\alpha}$ is the net force of all particles on the $i$-th particle in the $\alpha$-direction, and $r_{i}^{\beta}$ is the coordinate of the $i$-th particle in the $\beta$-direction. For the convenience of discussion below, we separates the kinetic stress $S_{0}^{\alpha\beta}$ that depends on velocities, and the excess stress $S_{\rm ex}^{\alpha\beta}$ that depends on the forces (the potential energy). Because the considered fluid is isotropic, we assume, $S^{xy}=S^{yz}=S^{xz}$.
S2.2.1 Kinetic theory
Using Eqs. (S19) and (S21), we can expand $G(t)$ into four terms,
$\begin{split}G(t)&=\left\langle S_{0}^{\alpha\beta}(0)S_{0}^{\alpha\beta}(t)\right\rangle+\left\langle S_{\rm ex}^{\alpha\beta}(0)S_{\rm ex}^{\alpha\beta}(t)\right\rangle+\left\langle S_{0}^{\alpha\beta}(0)S_{\rm ex}^{\alpha\beta}(t)\right\rangle+\left\langle S_{\rm ex}^{\alpha\beta}(0)S_{0}^{\alpha\beta}(t)\right\rangle\\ &\approx\left\langle S_{0}^{\alpha\beta}(0)S_{0}^{\alpha\beta}(t)\right\rangle+\left\langle S_{\rm ex}^{\alpha\beta}(0)S_{\rm ex}^{\alpha\beta}(t)\right\rangle\\ &\equiv G_{0}(t)+G_{\rm ex}(t).\end{split}$ (S22)
The third and fourth terms in the first line are neglected, assuming that the kinetic and potential parts are uncorrelated. In the kinetic theory, the excess part $G_{\rm ex}(t)$ is also ignored. In this case, $\eta$ is determined solely by velocity auto-correlations. Under similar approximations (e.g., the dynamics are memoryless), one can derive Chapman and Cowling (1990),
$\eta_{0}=\frac{5}{16\rho\sigma_{0}^{2}}\left(\frac{mk_{\rm B}T}{\pi}\right)^{1/2},$ (S23)
which has a simple relation with $D_{0}$, $\eta_{0}\sim D_{0}\rho$.
S2.2.2 Lucks-Moser theory
Lucks and Moser have also computed the excess viscosity by considering an exponential memory function Lucas and Moser (1979b). For the excess part, the stress auto-correlation function is,
$G_{\rm ex}(t)\equiv\left\langle S_{\rm ex}^{\alpha\beta}(0)S_{\rm ex}^{\alpha\beta}(t)\right\rangle,$ (S24)
which satisfies,
$\dot{G}_{\rm ex}(t)=-\int_{0}^{t}M_{G}(t-\tau)G_{\rm ex}(\tau)d\tau.$ (S25)
Using an exponential ansatz,
$M_{G}(t)=ae^{-bt},$ (S26)
where
$a=\frac{\left\langle\dot{S}^{\alpha\beta}(0)\dot{S}^{\alpha\beta}(0)\right\rangle}{\left\langle{S}^{\alpha\beta}(0){S}^{\alpha\beta}(0)\right\rangle},$ (S27)
$b=\left[\frac{\left\langle\ddot{S}^{\alpha\beta}(0)\ddot{S}^{\alpha\beta}(0)\right\rangle}{\left\langle\dot{S}^{\alpha\beta}(0)\dot{S}^{\alpha\beta}(0)\right\rangle}-\frac{\left\langle\dot{S}^{\alpha\beta}(0)\dot{S}^{\alpha\beta}(0)\right\rangle}{\left\langle{S}^{\alpha\beta}(0){S}^{\alpha\beta}(0)\right\rangle}\right]^{1/2},$ (S28)
the excess viscosity can be evaluated,
$\eta_{\rm ex}=\frac{b}{Vk_{\rm B}Ta},$ (S29)
and the total viscosity is,
$\eta=\eta_{0}+\eta_{\rm ex}.$ (S30)
Note that the $\eta_{\rm ex}$ depends $g(r)$ through the following expressions,
$\frac{\left\langle{S}^{\alpha\beta}(0){S}^{\alpha\beta}(0)\right\rangle}{V(kT)^{2}\left(1/\sigma_{0}^{3}\right)}=n^{*}+\frac{16\pi}{15}\frac{n^{*2}}{T^{*}}\int_{0}^{\infty}g_{2}\left(r_{12}^{*}\right)\left[5W_{12}^{\prime}+2W_{12}^{\prime\prime}r_{12}^{*2}\right]r_{12}^{*4}dr_{12}^{*},$ (S31)
$\begin{aligned} &\frac{\left\langle\dot{S}^{\alpha\beta}(0)\dot{S}^{\alpha\beta}(0)\right\rangle}{V(kT)^{2}\left(\epsilon/m\sigma_{0}^{5}\right)}=\frac{1024\pi}{15}\frac{n^{*2}}{T^{*}}\int_{0}^{\infty}g_{2}\left(r_{12}^{*}\right)\left[10W_{12}^{\prime}+4W_{12}^{\prime}W_{12}^{\prime\prime}r_{12}^{*2}+W_{12}^{\prime\prime}r_{12}^{*4}\right]r_{12}^{*4}dr_{12}^{*},\end{aligned}$ (S32)
$\begin{aligned} \frac{\left\langle\ddot{S}^{\alpha\beta}(0)\ddot{S}^{\alpha\beta}(0)\right\rangle}{V(kT)^{2}\left(\epsilon^{2}/m^{2}\sigma_{0}^{7}\right)}=&\frac{2048\pi}{15}\frac{n^{*2}}{T^{*}}\int_{0}^{\infty}g_{2}\left(r_{12}{}^{*}\right)\left[T^{*}\left(90W_{12}^{\prime}+180W_{12}^{\prime}W_{12}^{\prime\prime}r_{12}{}^{*2}\right.\right.+24W_{12}^{\prime}W^{\prime\prime\prime}{}_{12}r_{12}{}^{*4}\\ &+231W^{\prime\prime}{}_{12}{}^{2}r_{12}{}^{*4}+84W^{\prime\prime}{}_{12}W^{\prime\prime\prime}{}_{12}r_{12}{}^{*6}\left.+12W^{\prime\prime\prime}{}_{12}{}^{2}r_{12}{}^{*8}\right)+4\left(20W_{12}^{\prime}+24W_{12}^{\prime}W^{\prime\prime}{}_{12}r_{12}{}^{*2}\right.\\ &\left.\left.+18W_{12}^{\prime}W^{\prime\prime}{}_{12}{}^{2}r_{12}{}^{*4}+4W^{\prime\prime}{}_{12}{}^{3}r_{12}{}^{*6}\right)r_{12}{}^{*2}\right]r_{12}{}^{*2}dr_{12}{}^{*},\end{aligned}$ (S33)
where
$\begin{aligned} W_{ij}^{\prime\prime\prime}&=d^{3}W_{ij}/\left(dr_{ij}{}^{*2}\right)^{3}.\end{aligned}$ (S34)
S3 Supplementary analysis of velocity auto-correlation functions (VACFs)
In Fig.S3, we plot a complete set of VACFs measured by molecular dynamics (MD) simulations at $T=195$ K and different pressures.
Next, we detail the fitting procedure used in Fig. 4a. The VACF is computed for all simulated pressures and fitted over a time window chosen to ensure convergence while minimizing numerical noise from the long-time tail. Because the convergence rate of the VACF strongly depends on pressure—being slower at low pressures and faster at high pressures—the fitting window for the stretched-exponential form is adjusted accordingly. For instance, at $5$ bar the VACF is evaluated over $100$ ps and fitted in ten consecutive $5$ ps segments starting from $5$ ps. As pressure increases and correlations decay more rapidly, progressively shorter fitting windows are used: $80$ ps at $10$ bar, $60$ ps at $15$ bar, $50$ ps at 20202525 bar, and 4466 ps for pressures above $200$ bar. For the highest-pressure states ($\geq 400$ bar), the VACF converges within $4$ ps, and the fitting is performed in ten consecutive $0.1$ ps segments starting from $0.8$ ps.
Figure S3: Velocity autocorrelation functions. Velocity auto-correlation functions $Z(t)$ for argon are obtained from MD simulations at P=5,10,15,20,25,35,45,55,60,75,90,100,110,120,135,150,200,210,320,400,500,600,700,800,900,1000,1100,1200,1300barP=5,10,15,20,25,35,45,55,60,75,90,100,110,120,135,150,200,210,320,400,500,600,700,800,900,1000,1100,1200,1300\ $\text${bar} (from right to left) and T=195KT=195\ $\text${K}.
S4 Calculation of $S(q)$ and Ornstein-Zernike Correlation Length from MD data.
The static structure factor $S(q)$ is computed directly from molecular dynamics trajectories. For each configuration snapshot containing $N=4000$ argon atoms, $S(\mathbf{q})$ is calculated as:
$S(\mathbf{q})=\frac{1}{N}\left|\sum_{j=1}^{N}e^{-i\mathbf{q}\cdot\mathbf{r}_{j}}\right|^{2},$ (S35)
where $\mathbf{r}_{j}$ are atomic positions and $\mathbf{q}$ are wave vectors compatible with the periodic boundary conditions. The wave vectors are grouped into spherical shells of width $\Delta q=0.05$ Å-1, and $S(\mathbf{q})$ values within each shell are averaged to obtain $S(q)$. For each pressure, $S(q)$ is evaluated over 500 ps and averaged over 500,000 configurations.
The Ornstein-Zernike (OZ) correlation length is obtained by fitting the low-$q$ region of $S(q)$ to the OZ form:
$S(q)=\frac{S_{0}}{1+\xi_{\mathrm{OZ}}^{2}q^{2}}.$ (S36)
Inverting this relation, one gets
$\frac{1}{S(q)}=\frac{1}{S_{0}}+\frac{\xi_{\mathrm{OZ}}^{2}}{S_{0}}q^{2},$ (S37)
from which $\xi_{\mathrm{OZ}}$ is determined via linear regression. Uncertainties are estimated from the fitting errors.
Figure S4: Ornstein-Zernike correlation length. $\xi_{\mathrm{OZ}}$ as a function of pressure at $T=195$ K, showing a maximum near the Widom line between the $L^{+}$ and $L^{-}$ lines. Error bars represent fitting uncertainties.
The resulting $\xi_{\mathrm{OZ}}$ as a function of pressure is shown in Fig. S4. $\xi_{\mathrm{OZ}}$ reaches its maximum near the Widom line, which is located roughly in between the $L^{\pm}$ lines.
S5 Identification of the Widom delta
Following the approach of Yoon et al. Jun et al. (2018), we employ a probabilistic classification algorithm to characterize the gas-like and liquid-like domains in the simulated argon systems.
For each configuration from MD data, we perform the Voronoi tessellation. The Voronoi cell of atom $i$ is defined as the set of points closer to atom $i$ than to any other atom. The local density $\rho_{i}^{\text{local}}$ is then calculated as the inverse of the Voronoi cell volume:
$\rho_{i}^{\text{local}}=\frac{1}{V_{i}^{\text{Voronoi}}},$ (S38)
where $V_{i}^{\text{Voronoi}}$ is the volume of the Voronoi polyhedron associated with atom $i$.
To account for spatial correlations and reduce sensitivity to local fluctuations, we compute the mean local density $\bar{\rho}_{i}$ for each atom $i$ by averaging over its Voronoi neighbors:
$\bar{\rho}_{i}=\frac{1}{N_{\text{neigh}}+1}\left(\rho_{i}^{\text{local}}+\sum_{j\in\mathcal{N}(i)}\rho_{j}^{\text{local}}\right),$ (S39)
where $\mathcal{N}(i)$ denotes the set of Voronoi neighbors of atom $i$ (atoms sharing a Voronoi face), and $N_{\text{neigh}}=|\mathcal{N}(i)|$ is the number of such neighbors.
The classification of atoms into gas-like and liquid-like states follows a probabilistic algorithm inspired by the work of Yoon et al. Jun et al. (2018):
  1. 1.
    Random processing order: Atoms are processed in random order. When an atom is selected, the mean local density of that atom and all its Voronoi neighbors is computed. This neighborhood averaging reduces sensitivity to local fluctuations.
  2. 2.
    Threshold comparison: For the current unlabeled atom $i$, its mean local density $\bar{\rho}_{i}$ is compared to a threshold value $\rho_{\text{th}}$. We use the critical density of argon $\rho_{c}$ as the threshold:
    $\text{label}=\begin{cases}1&(\text{liquid-like})\quad\text{if }\bar{\rho}_{i}\geq\rho_{c},\\ 0&(\text{gas-like})\quad\text{if }\bar{\rho}_{i}<\rho_{c}.\end{cases}$ (S40)
  3. 3.
    Simultaneous labeling: The chosen atom $i$ and all its unlabeled Voronoi neighbors are assigned the same label, reflecting the correlated nature of local environments.
  4. 4.
    Multiple trials: Steps 1–3 are repeated $N_{\text{trials}}=300$ times with different random processing orders to obtain statistically robust results.
  5. 5.
    Probability calculation: The liquid-like probability $P_{i}^{\text{liq}}$ for atom $i$ is calculated as the fraction of trials in which it is classified as liquid-like:
    $P_{i}^{\text{liq}}=\frac{1}{N_{\text{trials}}}\sum_{t=1}^{N_{\text{trials}}}L_{i}^{(t)},$ (S41)
    where $L_{i}^{(t)}=1$ if atom $i$ is labeled liquid-like in trial $t$, and $L_{i}^{(t)}=0$ otherwise.
For each pressure, we calculate the fractions of gas-like and liquid-like atoms as ensemble averages. The gas-like fraction $\Pi_{\text{gas}}$ is defined as:
$\Pi_{\text{gas}}=\frac{1}{N_{\text{atoms}}}\sum_{i=1}^{N_{\text{atoms}}}\mathcal{I}\left(P_{i}^{\text{liq}}\leq 0.5\right),$ (S42)
where $\mathcal{I}(\cdot)$ is the indicator function. Similarly, the liquid-like fraction $\Pi_{\text{liq}}$ is:
$\Pi_{\text{liq}}=1-\Pi_{\text{gas}}.$ (S43)
The fractions are averaged over 100 random configurations along the trajectories. Uncertainties in these fractions are estimated from the standard deviation across multiple simulation frames.
The resulting fractions of gas-like and liquid-like atoms as functions of pressure for simulated argon at two temperatures is shown in Fig.S5. The gas-like fraction $\Pi_{\text{gas}}$ approaches unity near the $L^{-}$ line, while the liquid-like fraction $\Pi_{\text{liquid}}$ approach unity much earlier than the $L^{+}$ line.
The Widom delta is defined from the inverse sigmoidal fit to $\Pi_{\text{gas}}$ by finding the densities (pressures) where the tangent line at $\Pi_{\text{gas}}=0.5$ intersects $\Pi_{\text{gas}}=1.0$ and $\Pi_{\text{gas}}=0.0$, respectively. We find that the Widom delta is within the intermediate transition regime bounded by $L^{\pm}$ lines.
Figure S5: Comparison between Widom delta and $L^{\pm}$ lines.Fractions of gas-like ($\Pi_{\text{gas}}$) and liquid-like ($\Pi_{\text{liquid}}$) atoms as functions of pressure for simulated argon at two temperatures: $\mathbf{(a)}$ $T=195\,\mathrm{K}$ and $\mathbf{(b)}$ $T=220\,\mathrm{K}$. The red dashed line indicates the Widom line.